### v1

# Understanding the Training and Generalization of Pretrained Transformer for Sequential Decision Making

Abstract
In this paper, we consider the supervised pretrained transformer for a class of sequential decision-making problems. The class of considered problems is a subset of the general formulation of reinforcement learning in that there is no transition probability matrix, and the class of problems covers bandits, dynamic pricing, and newsvendor problems as special cases. Such a structure enables the use of optimal actions/decisions in the pretraining phase, and the usage also provides new insights for the training and generalization of the pretrained transformer. We first note that the training of the transformer model can be viewed as a performative prediction problem, and the existing methods and theories largely ignore or cannot resolve the arisen out-of-distribution issue. We propose a natural solution that includes the transformer-generated action sequences in the training procedure, and it enjoys better properties both numerically and theoretically. The availability of the optimal actions in the considered tasks also allows us to analyze the properties of the pretrained transformer as an algorithm and explains why it may lack exploration and how this can be automatically resolved. Numerically, we categorize the advantages of the pretrained transformer over the structured algorithms such as UCB and Thompson sampling into three cases: (i) it better utilizes the prior knowledge in the pretraining data; (ii) it can elegantly handle the misspecification issue suffered by the structured algorithms; (iii) for short time horizon such as $T\leq 50$, it behaves more greedy and enjoys much better regret than the structured algorithms which are designed for asymptotic optimality.
**footnotetext: Equal contribution.footnotetext: Corresponding to: Hanzhao Wang, h.wang19@imperial.ac.uk
1 Introduction
In recent years, transformer-based models (Vaswani et al., 2017) have achieved great success in a wide range of tasks such as natural language processing (Radford et al., 2019; Bubeck et al., 2023), computer vision (Dosovitskiy et al., 2020), and also reinforcement learning (RL) (Chen et al., 2021). In particular, an offline RL problem can be framed as a sequence prediction problem (Ding et al., 2019) that predicts the optimal/near-optimal/human action based on the observed history. Chen et al. (2021); Janner et al. (2021) pioneer this approach with the Decision Transformer (DT), which leverages the auto-regressive nature of transformers to maximize the likelihood of trajectories in offline datasets. In this paper, we apply the idea of supervised pretrained transformer on a subclass of reinforcement learning problems, which we call sequential decision-making problems, that has no transition probability matrix compared to a general RL problem. This simplified problem structure brings two benefits: (i) the optimal action is either in closed form or efficiently solvable and thus we can use the optimal action as the prediction target in the training data and (ii) the setting is more aligned with the existing study (Xie et al., 2021; Garg et al., 2022) on the in-context learning ability of the transformer which gives more insight to the working machinery of transformer on decision-making tasks. Moreover, the class of sequential decision-making problems studied in this paper, covers stochastic and linear bandit problems (Lattimore and Szepesvári, 2020), dynamic pricing (Den Boer, 2015; Cohen et al., 2020), and newsvendor problem (Levi et al., 2015; Besbes and Muharremoglu, 2013) as special cases. Therefore, as a side product, our paper extends pretrained transformers to the application contexts of business, economics, and operations research which are less studied by the existing works.
To summarize, we work on the following aspects in this paper:
- The training procedure: In generating the training data for the pretrained transformer, it needs to generate both the actions in the history (which are the features) and the optimal/near-optimal action to predict (which are the target variable). We point out that when these two streams of actions are generated by different algorithms/methods, it will result in an out-of-distribution (OOD) issue largely ignored or unsolved by the existing literature. Zheng et al. (2022) show that when both streams are generated by one structured algorithm (say UCB or Thompson sampling), the pretraining of the transformer will result in an imitation learning procedure that imitates the structured algorithm at best. If one desires to obtain a better algorithm than the structured algorithm, like using the optimal action (Lee et al., 2024) as the target to predict in our case, this will cause the OOD issue (reflected by an $\Theta(\exp T)$ factor in the regret bound of Zheng et al. (2022)). To resolve this matter, we propose an algorithmic solution by injecting the transformer-generated action sequence into the pretraining data, and this also draws a connection with the problem of performative prediction (Perdomo et al., 2020).
- Transformer as decision-making algorithms: The transformer is trained to minimize the empirical risk as the standard supervised learning, and hence the pretrained transformer will approximate the Bayes-optimal predictor for the supervised loss as best. Yet, when used for decision-making problems, the transformer works as a decision-making algorithm. We investigate the gap between the prediction and the decision-making by interpreting the transformer from a Bayesian posterior perspective (Xie et al., 2021; Zhang et al., 2023b; Bai et al., 2024), and in particular, for the continuous action space. The investigation explains (i) why the transformer model exhibits a lack of exploration and (ii) how it achieves a paramount or better performance than the structured algorithms.
- When the transformer is better: While there have been many structured algorithms for each of the decision-making problems that the transformer model is applied to, the question is whether and when the transformer can achieve a better performance than these algorithms. We use numerical experiments to attribute the advantage of the pretrained transformer into three scenarios: (i) it better utilizes the prior knowledge in the pretraining data; (ii) it can elegantly handle the misspecification issue suffered by the structured algorithms; (iii) for short time horizon such as $T\leq 50$ or $100$, it behaves more greedy and enjoys much better regret than the structured algorithms which are designed for asymptotic optimality.
We defer more discussions on the related works to Section A.
2 Problem Setup
In this section, we introduce a general formulation for sequential decision-making and describe the setup of supervised pretraining. An agent (decision maker) makes decisions in a sequential environment. The full procedure is described by the sequence
$(X_{1},a_{1},o_{1},X_{2},a_{2},o_{2},\ldots,X_{T},a_{T},o_{T}).$
Here $T$ is the horizon. At each time $t=1,\ldots,T$, the agent first observes a context vector $X_{t}\in\mathcal{X}\subset\mathbb{R}^{d}$ and takes the action/decision $a_{t}\in\mathcal{A}$, and then observes an outcome $o_{t}\in\mathcal{O}\subset\mathbb{R}^{k}$. Define the history
$H_{t}\coloneqq(X_{1},a_{1},o_{1},\ldots,X_{t-1},a_{t-1},o_{t-1},X_{t}),$
and the agent’s decision $a_{t}$ is based on $H_{t}$, i.e., non-anticipatory. In this light, we can model the decision as a function of $H_{t}$,
$a_{t}=\texttt{TF}_{\theta}(H_{t})$
where $\theta$ encapsulates all the parameters of the function. That is, the agent learns from the past interactions with the underlying environment – tuples of $(X_{t},a_{t},o_{t})$’s to optimize future decisions, and this process that transfers the past knowledge to the future is described by the function $\texttt{TF}_{\theta}$. The function $\texttt{TF}_{\theta}$ can take as inputs a variable-length sequence ($H_{t}$ can be of variable length), which is the case for the generative pretrained transformers (Radford et al., 2019).
2.1 Environment and performance metrics
The function $\texttt{TF}_{\theta}$ describes how the actions $a_{t}$’s are generated. Now we describe how $X_{t}$’s and $o_{t}$’s are generated. Assume the context $X_{t}$’s are i.i.d. and $o_{t}$ follows a conditional distribution given $X_{t}$ and $a_{t}$
[…]
where $\gamma$ encapsulates all the parameters for these two probability distributions. In particular, we emphasize that the parameter $\gamma$ is unknown to the agent.
At each time $t$, the agent collects a random reward $R_{t}=R(X_{t},a_{t})$ which depends on $X_{t}$, $a_{t},$ and also some possible exogenous randomness. The observation $o_{t}$ includes $R_{t}$ as its first coordinate. The expected reward is defined by $r(X_{t},a_{t})=\mathbb{E}[R(X_{t},a_{t})|X_{t},a_{t}]$. Specifically, this reward function $r(\cdot,\cdot)$ is unknown to the agent and it also depends on the underlying environment $\gamma$.
The performance of the agent, namely, the decision function $\texttt{TF}_{\theta}$, is usually measured by the notion of regret, which is defined by
$\text{Regret}(\texttt{TF}_{\theta};\gamma)\coloneqq\mathbb{E}\left[\sum_{t=1}^{T}r(X_{t},a_{t}^{*})-r(X_{t},a_{t})\right]$ (2)
where the action $a_{t}^{*}$ is the optimal action that maximizes the reward
$a_{t}^{*}\coloneqq\operatorname*{arg\,max}_{a\in\mathcal{A}}\ r(X_{t},a)$ (3)
and the optimization assumes the knowledge of the underlying environment, i.e., the parameter $\gamma$. In the regret definition (2), the expectation is taken with respect to the underlying probability distribution (1) and possible randomness in the agent’s decision function $\texttt{TF}_{\theta}$. We use the arguments $\texttt{TF}_{\theta}$ and $\gamma$ to reflect the dependency of the regret on the decision function $\texttt{TF}_{\theta}$ and the environment $\gamma.$ Also, we note that the benchmark is defined as a dynamic oracle where $a_{t}^{*}$ maybe different over time. Regret is not a perfect performance measure for the problem, and we also do not believe the algorithm design for sequential decision-making problems should purely aim for getting a better regret bound. Despite this, it serves as a good monitor for the effectiveness of the underlying learning and the optimization procedure.
2.2 Supervised pretraining
In the classic study of these sequential decision-making problems, the common paradigm is to assume some structure for the underlying environment and/or the reward function, and to design algorithms accordingly. These algorithms are often known as online or reinforcement learning algorithms and are usually hard to combine with prior knowledge such as pretraining data. Comparatively, the supervised pretraining formulates the decision-making problem as a supervised learning prediction task that predicts the optimal or a near-optimal action, and thus it can utilize the vast availability of pretraining data that can be generated from simulated environments.
Specifically, the supervised pretraining first constructs a pretraining dataset
$\mathcal{D}_{\text{PT}}\coloneqq\left\{\left(H_{1}^{(i)},a_{1}^{(i)*}\right),\left(H_{2}^{(i)},a_{2}^{(i)*}\right),\ldots,\left(H_{T}^{(i)},a_{T}^{(i)*}\right)\right\}_{i=1}^{n}.$
In particular, we assume the parameter $\gamma$ (see (1)) that governs the generation of $X_{t}$ and $o_{t}$ are generated from some environment distribution $\mathcal{P}_{\gamma}$. Then, we first generate $n$ environments denoted by $\gamma_{1},\ldots,\gamma_{n}.$ Then there are two important aspects with regard to the generation of $\mathcal{D}_{\text{PT}}$:
- Action $a_{t}^{(i)*}$: For this paper, we generate $a_{t}^{(i)*}$ as the optimal action specified by (3) and this requires the knowledge of the underlying environment $\gamma_{i}.$ A common property of all these sequential decision-making problems is that the optimal action can be easily or often analytically computed. Other option is also possible, say, $a_{t}^{(i)*}$ can be generated based on some expert algorithms such as Thompson sampling or UCB algorithms. In this case, the action $a_{t}^{(i)*}=\texttt{Alg}(H_{t}^{(i)})$ where $\texttt{Alg}(\cdot)$ represents the algorithm that maps from $H_{t}^{(i)}$ to the action without utilizing knowledge of $\gamma_{i}.$
- History $H_{t}^{(i)}=\left(X_{1}^{(i)},a_{1}^{(i)},o_{1}^{(i)},\ldots,X_{t-1}^{(i)},a_{t-1}^{(i)},o_{t-1}^{(i)},X_{t}^{(i)}\right).$ Note that $X_{t}^{(i)}$’s and $o_{t}^{(i)}$’s are generated based on the parameter $\gamma_{i}$ following (1). The action $a_{t}^{(i)}$ is generated based on some pre-specified decision function $f$ in a recursive manner:
$a_{t}^{(i)}=f\left(H_{t}^{(i)}\right),\ \ H_{t+1}^{(i)}=\left(H_{t}^{(i)},a_{t}^{(i)},o_{t}^{(i)},X_{t+1}^{(i)}\right)$ (4)
for $t=1,\ldots,T.$
Formally, we define a distribution $\mathcal{P}_{\gamma,f}$ that generates $H_{t}$ and $a_{t}^{*}$ as
$X_{\tau}\sim\mathbb{P}_{\gamma}(\cdot),\ a_{\tau}=f(H_{\tau}),\ o_{\tau}\sim\mathbb{P}_{\gamma}\left(\cdot|X_{\tau},a_{\tau}\right),\ H_{\tau+1}=\left(H_{\tau},a_{\tau},o_{\tau},X_{\tau+1}\right).$ (5)
for $\tau=1,\ldots,t-1$ and $a_{t}^{*}$ is specified by (3). The parameter $\gamma$ and the function $f$ jointly parameterize the distribution $\mathcal{P}_{\gamma,f}$. The parameter $\gamma$ governs the generation of $X_{t}$ and $o_{t}$ just as in (1). The decision function $f$ maps from the history to the action and it governs the generation of $a_{\tau}$’s in the sequence $H_{t}$. In this way, the pretraining data is generated with the following flow:
$\mathcal{P}_{\gamma}\rightarrow\gamma_{i}\rightarrow\mathcal{P}_{\gamma_{i},f}\rightarrow\left\{\left(H_{1}^{(i)},a_{1}^{(i)*}\right),\ldots,\left(H_{T}^{(i)},a_{T}^{(i)*}\right)\right\}$
where $f$ is a prespecified decision function used to generate the data.
Based on the dataset $\mathcal{D}_{\text{PT}}$, the supervised pretraining refers to the learning of the decision function $\texttt{TF}_{\theta}$ through minimizing the following empirical loss
$\hat{\theta}\coloneqq\operatorname*{arg\,min}_{\theta}\frac{1}{nT}\sum_{i=1}^{n}\sum_{t=1}^{T}l\left(\texttt{TF}_{\theta}\left(H_{t}^{(i)}\right),a_{t}^{(i)*}\right)$ (6)
where the loss function $l(\cdot,\cdot):\mathcal{A}\times\mathcal{A}\rightarrow\mathbb{R}$ specifies the prediction loss. Thus the key idea of supervised pretraining is to formulate the sequential decision-making task as a sequence modeling task that predicts the optimal action $a_{t}^{(i)*}$ given the history $H_{t}^{(i)}$.
In the test phase, a new environment parameter $\gamma$ is generated from $\mathcal{P}_{\gamma}$ and we apply the learned decision function $\texttt{TF}_{\hat{\theta}}$. Specifically, the procedure is described by
$X_{\tau}\sim\mathbb{P}_{\gamma}(\cdot),\ a_{\tau}=\texttt{TF}_{{\hat{\theta}}}(H_{\tau}),\ o_{\tau}\sim\mathbb{P}_{\gamma}\left(\cdot|X_{\tau},a_{\tau}\right),\ H_{\tau+1}=\left(H_{\tau},a_{\tau},o_{\tau},X_{\tau+1}\right).$ (7)
The only difference between the test dynamics (7) and the training dynamics (5) is that the action $a_{t}$ (or $a_{\tau}$) is generated by $\texttt{TF}_{{\hat{\theta}}}$ instead of the pre-specified decision function $f$. In this way, we can write $(H_{t},a_{t}^{*})\sim\mathcal{P}_{\gamma,\texttt{TF}_{\theta}}$ for the test phase. Accordingly, we can define the expected loss of the decision function $\texttt{TF}_{\theta}$ under the environment specified by $\gamma$ as
$L\left(\texttt{TF}_{\theta};\gamma\right)\coloneqq\mathbb{E}_{(H_{t},a_{t}^{*})\sim\mathcal{P}_{\gamma,\texttt{TF}_{\theta}}}\left[\sum_{t=1}^{T}l\left(\texttt{TF}_{\theta}\left(H_{t}\right),a_{t}^{*}\right)\right]$ (8)
where the expectation is taken with respect to $X_{t}$, $o_{t}$ and the possible randomness in $a_{t}.$ As before, $a_{t}^{*}$ are the optimal action defined by (3) and obtained from the knowledge of $\gamma$.
We defer discussions on how various sequential decision-making problems can be formulated under the general setup above and how the pretraining loss is related to the regret to Section B.
3 Pretraining and Generalization
We first point out a subtle point of the supervised pretraining setting presented in the previous section. Unlike supervised learning, the expected loss (8) may not be equal to the expectation of the empirical loss (6). As noted earlier, this is because in the sequence $H_{t}$, the action $a_{t}=\texttt{TF}_{{\hat{\theta}}}(H_{t})$ as in (7) during the test phase as opposed to that $a_{t}=f(H_{t})$ as in (5) for the pertraining data $\mathcal{D}_{\text{PT}}$. Taking expectation with respect to $\gamma$ in (8), we can define the expected loss as
$L(\texttt{TF}_{\theta})\coloneqq\mathbb{E}_{\gamma\sim\mathcal{P}_{\gamma}}\left[L(\texttt{TF}_{\theta};\gamma)\right].$
However, the expectation of the empirical loss (6) is
$L_{f}(\texttt{TF}_{\theta})\coloneqq\mathbb{E}_{\gamma\sim\mathcal{P}_{\gamma}}\left[L_{f}(\texttt{TF}_{\theta};\gamma)\right]$ (9)
where
$L_{f}(\texttt{TF}_{\theta};\gamma)\coloneqq\mathbb{E}_{(H_{t},a_{t}^{*})\sim\mathcal{P}_{\gamma,f}}\left[\sum_{t=1}^{T}l\left(\texttt{TF}_{\theta}\left(H_{t}\right),a_{t}^{*}\right)\right].$
Here $f$ is the pre-specified decision function used in generating the pretraining data $\mathcal{D}_{\text{PT}}$. We note that the discrepancy between $L(\texttt{TF}_{\theta})$ and $L_{f}(\texttt{TF}_{\theta})$ is caused by the difference in the generations of the samples $(H_{t},a_{t}^{*})$, namely, $\mathcal{P}_{\gamma,\texttt{TF}_{\theta}}$ versus $\mathcal{P}_{\gamma,f}$; which essentially reduces to how the actions $a_{t}$’s are generated in $H_{t}$. In the training phase, this is generated based on $f$, while measuring the test loss requires that the actions to be generated are based on the transformer $\texttt{TF}_{\theta}.$ This discrepancy is somewhat inevitable because when generating the training data, there is no way we can know the final parameter $\hat{\theta}.$ Consequently, there is no direct guarantee on the loss for $L(\texttt{TF}_{\hat{\theta}};\gamma)$ or $L(\texttt{TF}_{\hat{\theta}})$ with $\hat{\theta}$ being the final learned parameter. This out-of-distribution (OOD) issue motivates the design of our algorithm.
Algorithm 1 Supervised pretraining transformers for sequential decision making
1:Iterations $M$, early training phase $M_{0}\in[1,M]$, number of training sequences per iteration $n$, ratio $\kappa\in[0,1]$, pretraining decision function $f$, prior distribution $\mathcal{P}_{\gamma}$, initial parameter $\theta_{0}$
2:Initialize $\theta_{1}=\theta_{0}$
3:for $m=1,\ldots,M_{0}$ do
4:    %Early training phase
5:     Sample $\gamma_{1},\gamma_{2},\ldots,\gamma_{n}$ from $\mathcal{P}_{\gamma}$ and $S_{i}=\left\{\left(H_{1}^{(i)},a_{1}^{(i)*}\right),\ldots,\left(H_{T}^{(i)},a_{T}^{(i)*}\right)\right\}$ from $\mathcal{P}_{\gamma_{i},f}$
6:     Optimize over the generated dataset $\mathcal{D}_{m}=\{S_{i}\}_{i=1}^{n}$ and obtain the updated parameter $\theta_{m+1}$
7:end for
8:for $m=M_{0}+1,\ldots,M$ do
9:    %Mixed training phase
10:     Sample $\gamma_{1},\gamma_{2},\ldots,\gamma_{n}$ from $\mathcal{P}_{\gamma}$
11:     For $i=1,\ldots,\kappa n$, sample $S_{i}=\left\{\left(H_{1}^{(i)},a_{1}^{(i)*}\right),\ldots,\left(H_{T}^{(i)},a_{T}^{(i)*}\right)\right\}$ from $\mathcal{P}_{\gamma_{i},f}$
12:     For $i=\kappa n,\ldots,n$, sample $S_{i}=\left\{\left(H_{1}^{(i)},a_{1}^{(i)*}\right),\ldots,\left(H_{T}^{(i)},a_{T}^{(i)*}\right)\right\}$ from $\mathcal{P}_{\gamma_{i},\texttt{TF}_{\theta_{m}}}$
13:     Optimize over the generated dataset $\mathcal{D}_{m}=\{S_{i}\}_{i=1}^{n}$ and obtain the updated parameter $\theta_{m+1}$
14:end for
15:Return: $\hat{\theta}=\theta_{M+1}$
Algorithm 1 describes our algorithm to resolve the above-mentioned OOD issue between training and test. It consists of two phases, an early training phase and a mixed training phase. For the early training phase, the data are generated from the distribution $\mathcal{P}_{\gamma_{i},f}$, while for the mixed training phase, the data are generated from both $\mathcal{P}_{\gamma_{i},f}$ and $\mathcal{P}_{\gamma_{i},\texttt{TF}_{\theta_{t}}}$ with a ratio controlled by $\kappa$. The decision function $f$ can be chosen as a uniform distribution over the action space as in the literature (Chen et al., 2021; Zhang et al., 2023a; Lee et al., 2024), or other more complicated options (details on the $f$ used in our experiments are provided in Appendix E.1.2). When there is no mixed phase, i.e., $M_{0}=M$, the algorithm recovers that of (Lee et al., 2024). Intuitively, the benefits of involving the data generated from $\mathcal{P}_{\gamma_{i},\texttt{TF}_{\theta_{t}}}$ is to make the sequences $H_{t}^{(i)}$’s in the training data closer to the ones generated from the transformer. For the mixed training phase, a proportion of the data is still generated from the original $\mathcal{P}_{\gamma_{i},f}$ because the generation from $\mathcal{P}_{\gamma_{i},\texttt{TF}_{\theta_{t}}}$ costs relatively more time. For our numerical experiments, we choose $\kappa=1/3.$
(a) Training dynamics
(b) t-SNE of the generated $H_{30}$’s
Figure 1: (a) Training dynamics. Orange: $M_{0}=M=130.$ Blue: $M_{0}=50$ and $M=130$. It shows the effectiveness of injecting/mixing the transformer-generated sequence into the training procedure. (b) A visualization of the $H_{t}$ with $a_{\tau}$’s in $H_{t}$ generated from various $\texttt{TF}_{\theta_{m}}.$ For each $\texttt{TF}_{\theta_{m}}$, we generate 30 sequences. The decision function $\texttt{Alg}^{*}$ is defined in the next section. We observe (i) there is a shift over time in terms of the transformer-generated action sequence, and thus the training should adaptively focus more on the recently generated sequence like the design in Algorithm 1; (ii) the action sequence gradually gets closer to the optimal decision function $\texttt{Alg}^{*}.$ The experiment setups are deferred to Appendix E.4.
[…]
Algorithm 1 from an optimization perspective can be viewed as an iterative procedure to optimize
$\mathbb{E}_{\gamma\sim\mathcal{P}_{\gamma}}\left[\mathbb{E}_{(H_{t},a_{t}^{*})\sim\kappa\mathcal{P}_{\gamma,f}+(1-\kappa)\mathcal{P}_{\gamma,\texttt{TF}_{\theta}}}\left[\sum_{t=1}^{T}l\left(\texttt{TF}_{\theta}\left(H_{t}\right),a_{t}^{*}\right)\right]\right]$
where both the input data pair $(H_{t},a_{t}^{*})$ and the prediction function involve the parameter $\theta$. This falls into the paradigm of performative prediction (Perdomo et al., 2020; Izzo et al., 2021; Mendler-Dünner et al., 2020) where the prediction model may affect the data generated to be predicted. In the literature of performative prediction, a critical matter is the instability issue which may cause the parameter $\theta_{m}$ to oscillate and not converge. Perdomo et al. (2020) shows that the matter can be solved with strong conditions such as smoothness and strong convexity on the objective function. However, we do not encounter this instability in our numerical experiment, and we make an argument as the following claim. That is, when the underlying function class, such as $\{\texttt{TF}_{\theta}:\theta\in\Theta\}$ is rich enough, one does not need to worry about such instability.
Claim 3.1.
The instability of a performative prediction algorithm and the resultant non-convergence does not happen when the underlying prediction function is rich enough.
The claim will be further reinforced by the optimal decision function $\texttt{Alg}^{*}$ and Proposition 4.1 in the next section. We also defer more discussions on the claim to Section D.1. While it brings a peace of mind from the optimization perspective, the following proposition justifies the mixed training procedure from a statistical generalization perspective.
Proposition 3.2.
Suppose $\hat{\theta}$ is determined by (6) where $\kappa n$ data sequences are from $\mathcal{P}_{\gamma,f}$ and $(1-\kappa)n$ data sequences are from $\mathcal{P}_{\gamma,\texttt{TF}_{\tilde{\theta}}}$ for some parameter $\tilde{\theta}$. For a Lipschitz and bounded loss $l$, the following generalization bound holds with probability at least $1-\delta$,
$L(\texttt{TF}_{\hat{\theta}})$ $\leq\hat{L}(\texttt{TF}_{\hat{\theta}})+\sqrt{\frac{\mathrm{Complex}(\{\texttt{TF}_{\theta}:\theta\in\Theta\})}{nT}}$
$+O\left(\kappa T\cdot\mathbb{E}_{\gamma\sim\mathcal{P}_{\gamma}}\left[W_{1}\left(\mathcal{P}_{\gamma,f},\mathcal{P}_{\gamma,\texttt{TF}_{\hat{\theta}}}\right)\right]+(1-\kappa)T\cdot\mathbb{E}_{\gamma\sim\mathcal{P}_{\gamma}}\left[W_{1}\left(\mathcal{P}_{\gamma,\texttt{TF}_{\tilde{\theta}}},\mathcal{P}_{\gamma,\texttt{TF}_{\hat{\theta}}}\right)\right]+\sqrt{\frac{\log\left(\frac{1}{\delta}\right)}{nT}}\right)$
where $\mathrm{Complex}(\cdot)$ denotes some complexity measure and $W_{1}(\cdot,\cdot)$ is the Wasserstein-1 distance. The notation $O(\cdot)$ omits the boundedness parameters.
The proof of Proposition 3.2 is not technical and follows the standard generalization bound (Mohri et al., 2018) together with the OOD analysis (Ben-David et al., 2010). We hope to use the result to provide more insights into the training procedure. The parameter $\tilde{\theta}$ can be interpreted as close to $\hat{\theta}$ as in the design of our algorithm. This will result in the term related to the distance between $\mathcal{P}_{\gamma,\texttt{TF}_{\tilde{\theta}}}$ and $\mathcal{P}_{\gamma,\texttt{TF}_{\hat{\theta}}}$ small. At the two ends of the spectrum, when $\kappa=1$, it will result in a large discrepancy between the training and the test data and thus the right-hand-side will explode; when $\kappa=0$, this will result in the tightest bound but it will cause more computational cost in the training procedure.
4 Learned Decision Function as an Algorithm
In the previous section, we discuss the training and generalization aspects of the pretrained transformer. Now, we return to the perspective of sequential decision-making and discuss the properties of the decision function when it is used as an algorithm.
In the ideal case, we define the Bayes-optimal decision function as follow
$\texttt{Alg}^{*}(H)$ $\coloneqq\operatorname*{arg\,min}_{a\in\mathcal{A}}\mathbb{E}\left[l(a,a^{*}(H,\gamma))|H\right]$
$=\operatorname*{arg\,min}_{a\in\mathcal{A}}\int_{\gamma}l(a,a^{*}(H,\gamma))\mathbb{P}(H|\gamma)\mathrm{d}\mathcal{P_{\gamma}}$
where $\mathcal{P}_{\gamma}$ is the environment distribution that generates the environment parameter $\gamma$, the optimal action $a^{*}(H,\gamma)$ depends on both the history $H$ and the environment parameter $\gamma$ (as defined in (3)), and $\mathbb{P}(H|\gamma)$ is proportional to the likelihood of observing the history $H$ under the environment $\gamma$
$\mathbb{P}(H|\gamma)\propto\prod_{t=1}^{\tau}\mathbb{P}_{\gamma}(X_{t})\cdot\prod_{t=1}^{\tau-1}\mathbb{P}_{\gamma}(o_{t}|X_{t},a_{t})$
where $H=(X_{1},a_{1},o_{1},\ldots,X_{\tau-1},a_{\tau-1},o_{\tau-1},X_{\tau})$ for some $\tau\in\{1,\ldots,T\}$ and the distributions $\mathbb{P}_{\gamma}(\cdot)$ and $\mathbb{P}_{\gamma}(\cdot|X_{t},a_{t})$ are as in (1). In the definition, the integration is with respect to the environment distribution $\mathcal{P}_{\gamma}$, and each possible environment $\gamma$ is weighted with the likelihood $\mathbb{P}(H|\gamma)$.
The following proposition relates $\texttt{Alg}^{*}(H)$ with the pretraining objective which justifies why it is called the optimal decision function. The loss function $L_{f}(\texttt{Alg})$ is defined as (9); recall that the decision function $f$ is used to generate the actions in $H_{t}$ and the decision function Alg is used to predict $a_{t}^{*}.$
Proposition 4.1.
The following holds for any distribution $\mathcal{P}_{\gamma,f}$:
$\texttt{Alg}^{*}(\cdot)\in\operatorname*{arg\,min}_{\texttt{Alg}\in\mathcal{F}}\ L_{f}(\texttt{Alg})=\mathbb{E}_{\gamma\sim\mathcal{P}_{\gamma}}\left[\mathbb{E}_{(H_{t},a_{t}^{*})\sim\mathcal{P}_{\gamma,f}}\left[\sum_{t=1}^{T}l\left(\texttt{Alg}\left(H_{t}\right),a_{t}^{*}\right)\right]\right]$
where $\mathcal{F}$ is the family of all measurable functions (on a properly defined space that handles variable-length inputs). In particular, we can choose the decision function $f$ properly to recover the test loss $L(\texttt{TF}_{\theta})$ or the expected training loss $L_{f}(\texttt{TF}_{\theta})$ in Section 3.
(a) Dynamic pricing
(b) Newsvendor
Figure 2: The pretrained transformer $\texttt{TF}_{\hat{\theta}}$ well matches the optimal decision function $\texttt{Alg}^{*}$. For both figures, we plot decision trials for $\texttt{Alg}^{*}$ and $\texttt{TF}_{\hat{\theta}}$. The optimal actions change over time because $X_{t}$’s are different for different time $t$. The experiment setup is deferred to Appendix E.4.
Proposition 4.1 states that the function $\texttt{Alg}^{*}$ is one minimizer of the expected loss under any distribution $\mathcal{P}_{\gamma,f}.$ To see this, $\texttt{Alg}^{*}$ is defined in a pointwise manner for every possible $H.$ The probability $\mathcal{P}_{\gamma,f}$ defines a distribution over the space of $H$. The pointwise optimality of $\texttt{Alg}^{*}$ easily induces the optimality of $\texttt{Alg}^{*}$ for any $\mathcal{P}_{\gamma,f}$. Furthermore, this implies that $\texttt{Alg}^{*}$ is also the optimal solution to the expected test loss and the expected training loss. This means that if the transformer class of $\texttt{TF}_{\theta}$ is rich enough to cover $\texttt{Alg}^{*}$, and with an infinite amount of data, the supervised pretraining will result in $\texttt{Alg}^{*}$ at best. Thus we can interpret the property of the supervised pretrained transformer by $\texttt{Alg}^{*}$ as a proxy. Such a Bayes-optimal function and the Bayesian perspective have been discussed under various settings in the in-context learning literature (Xie et al., 2021; Zhang et al., 2023b; Jeon et al., 2024).
Example 4.2.
$\texttt{Alg}^{*}$ behaves as (a) posterior sampling under the cross-entropy loss; (b) posterior averaging under the squared loss; (c) posterior median under the absolute loss; where the posterior distribution is with respect to $\gamma$ and it is defined by $\mathcal{P}(\gamma|H)\propto\mathbb{P}(H|\gamma)\cdot\mathcal{P}_{\gamma}(\gamma)$.
We defer more details for Example 4.2 to Appendix D.4. Part (a) has been directly employed as Assumption 1 in Lee et al. (2024) for their theoretical developments. Indeed, part (a) corresponds to a discrete action space such as multi-armed bandits while part (b) and part (c) correspond to the continuous action space such as linear bandits, pricing, and newsvendor problems.
Proposition 4.3.
There exist a linear bandits problem instance and a dynamic pricing problem instance, i.e., an environment distribution $\mathcal{P}_{\gamma}$, such that the optimal decision function learned under the squared loss incurs a linear regret for every $\gamma$, i.e., $\mathrm{Regret}(\texttt{Alg}^{*};\gamma)=\Omega(T)$ for every $\gamma$.
Proposition 4.3 states a negative result that the optimal decision function $\texttt{Alg}^{*}$, albeit being optimal in a prediction sense (for predicting the optimal action), does not serve as a good algorithm for sequential decision-making problems. To see the intuition, the decision function is used recursively in the test phase (see (7)). While it gives the best possible prediction, it does not conduct exploration which is critical for the concentration of the posterior measure. This highlights one key difference between the sequential decision-making problem and the regression problem (Garg et al., 2022) when applying the pretrained transformer. For the regression problem, the in-context samples (in analogy context-action-observation tuples in $H_{t}$) are i.i.d. generated, and this results in a convergence/concentration of the posterior to the true model/environment $\gamma$ (Ghosal et al., 2000). Yet for sequential decision-making problems, the dynamics (7) induces a dependency throughout $H_{t}$, and it may result in a non-concentration behavior for the posterior. And this is the key for constructing the linear-regret example in Proposition 4.3 and the examples are also inspired from and share the same spirit as the discussions in (Harrison et al., 2012). Such lack of exploration does not happen for discrete action space (Lee et al., 2024); this is because for discrete action space, the output $\texttt{TF}_{\hat{\theta}}$ gives the posterior distribution as a distribution over the action space. The distribution output will lead to a randomized action and thus automatically perform exploration.
Proposition 4.4.
Consider a finite prior $\mathcal{P}_{\gamma}$ supported on $\Gamma=\{\gamma_{1},...,\gamma_{k}\}$. Suppose (i) The transformer $\texttt{TF}_{\hat{\theta}}$ satisfies $\mathbb{E}[r(X_{t},\texttt{Alg}^{*}(H_{t}))|H_{t}]-\mathbb{E}[r(X_{t},\texttt{TF}_{\hat{\theta}}(H_{t}))|H_{t}]\leq\epsilon$ for every $H_{t}$ where the expectation is taken with respect to the possible randomness in $\texttt{Alg}^{*}$ and $\texttt{TF}_{\hat{\theta}}$. (ii) Denote the conditional distribution of the observation $o_{t}$ given the history $H_{t}$ and the environment $\gamma_{i}$ as $\mathbb{P}_{\gamma_{i}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}}).$ Assume there exists $\Delta\geq 0$ such that the KL divergence
$\text{KL}\left(\mathbb{P}_{\gamma_{i}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})\|\mathbb{P}_{\gamma_{j}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})\right)\geq\Delta$
for any $\gamma_{i},\gamma_{j}\in\Gamma$ and $H_{t}$ and the likelihood ratio between $\mathbb{P}_{\gamma_{i}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})$ and $\mathbb{P}_{\gamma_{j}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})$ is sub-Gaussian when $\gamma_{i}$ is the sampled environment. Then we have
$\mathrm{Regret}(\texttt{TF}_{\theta};\gamma)\leq O\left(\epsilon T+\frac{\log k}{\Delta}\right),$
where the notation $O(\cdot)$ omits the boundedness parameters and the sub-Gaussian parameter.
There are simple remedies to the linear regret behavior of the optimal decision function, for example, using an $\epsilon$-greedy version of the transformer. Numerically, we do not observe such behavior for any experiment even without any further modification of the pretrained transformer. Proposition 4.4 provides a simple explanation for why one does not have to be too pessimistic. Intuitively, the pretrained model $\texttt{TF}_{\theta}$ can be viewed as the optimal decision function plus some random noise, and the random noise actually plays the role of exploration and helps the posterior to concentration. Proposition 4.4 involves two conditions. The first condition is that $\texttt{TF}_{\hat{\theta}}$ behaves closely to $\texttt{Alg}^{*}$ and the deviation between these two is captured by $\epsilon$. This condition can be justified by Figure 2. The second condition characterizes the KL divergence between the conditional distribution of the observation $o_{t}$ for any two environments. Note that this KL divergence critically depends on $\texttt{TF}_{\theta}$ because the ability to distinguish between two environments depends on the intensity of exploration of the underlying algorithm. These two conditions work more like a stylized model to analyze the behavior of the transformer. It is a largely simplified model; say, the parameter $\Delta$ should have a dependency on $t$ for many online learning algorithms such as UCB or Thompson sampling. Nevertheless, the two conditions give us a perspective to understand how the transformer works as an online algorithm. In particular, the transformer deviates from $\texttt{Alg}^{*}$ by an amount of $\epsilon$ and this deviation accumulates over time into the first term in the regret bound. On the other hand, this deviation gives an exploration, and the second term captures the regret caused during the concentration of the posterior onto the true environment. In Appendix D.6.1, we provide some justifications for the two conditions, and also establish $\Delta=\Theta(\epsilon)$ for linear bandits and dynamic pricing. We also note that the finiteness of $\mathcal{P}_{\gamma}$ is not essential. The KL divergence condition can be modified for a continuous $\mathcal{P}_{\gamma}$ and combined with a covering argument to obtain the bound accordingly.
5 Numerical Experiments and Discussions
The following figure summarizes the experiments for our model against benchmarks for each task. We note that the prior environment distribution $\mathcal{P}_{\gamma}$ takes an infinite support which means the environment $\gamma$ in the test phase is different from the environment $\gamma_{i}$’s used for training with probability 1. This provides a fair and possibly more challenging setup to examine the ability of the transformer.
(a) Multi-armed bandits
(b) Linear bandits
(c) Dynamic pricing
(d) Newsvendor
Figure 3: The average out-of-sample regret of $\texttt{TF}_{\hat{\theta}}$ against benchmark algorithms (see details in Appendix E.3) calculated based 100 runs. The numbers in the legend bar are the final regret at $t=100$. The shaded area in the plots indicates the $90\%$ (empirical) confidence interval for the regrets. The prior distribution $\mathcal{P}_{\gamma}$ is continuous (infinitely many possible $\gamma$). The problem dimension: number of arms for MAB =20, dimension of linear bandits = 2, dimension of $X_{t}$ for pricing = 6, dimension of $X_{t}$ for newsvendor = 4.
We make the following observations on the numerical experiment. The transformer achieves a better performance than the structured algorithms for each task. This result is impressive to us in that the transformer is not replicating/imitating the performance of these benchmark algorithms but it discovers a new and more effective algorithm. This in fact shows the effectiveness of using the optimal actions $a_{t}^{*}$ as the target variable for training the transformer. Moreover, we attribute the advantage of the transformer against the benchmark algorithms to two aspects: prior knowledge from pretraining data and more greedy exploitation. First, the transformer well utilizes the pretraining data and can be viewed as tailored for the pretraining distribution. Comparatively, all the benchmark algorithms are cold-start. Second, the benchmark algorithms are all designed in ways to cater for asymptotic optimality and thus may sacrifice short-term rewards. Comparatively, the transformer behaves more greedy; just like $\texttt{Alg}^{*}$ is a decision function that does not explicitly encourage exploration. And such greedy behavior brings better reward when $T$ is small such as $T\leq 100.$ More numerical experiments are deferred to Section C.
Conclusion and limitations. In this paper, we study the pretrained transformer for sequential decision-making problems. Such problems enable using the optimal action as the target variable for training the transformer. We investigate the training and generation aspect of the transformer and identify an OOD issue that is largely ignored by the existing literature. Moreover, we interpret how the transformer works as an online algorithm through the lens of the Bayes-optimal decision function. The numerical experiments are encouraging in that the trained transformer learns a decision function better than all the benchmark algorithms. However, we point out that all the numerical experiments in this paper are conducted under small dimensions with $\text{dim}(X_{t})\leq 10.$ Such settings provide a testbed for us to understand the working mechanism of the trained transformer; but to make it more practically useful, scaling up its ability for high-dimensional settings is necessary. Our conjecture is that this will entails a longer training procedure as well as a potential larger model architecture.
Appendix A More Related Works
In this section, we discuss more related works which complement our discussions in Section 1. In particular, we elaborate the literature in three streams.
Understanding transformer and in-context learning.
A remarkable characteristic of pretrained transformers is their ability to perform in-context learning (ICL). Once pretrained on a vast corpus, transformers can solve new tasks with just a few examples, without updating their parameters. ICL has captured the attention of the theoretical machine learning community, leading to considerable efforts into understanding ICL from different theoretical perspectives [Xie et al., 2021, Akyürek et al., 2022, Von Oswald et al., 2023, Lin et al., 2023]. A stream of literature that is related to our work explains the transformer’s behavior as Bayesian inference. See Xie et al. [2021], Zhang et al. [2023b] for the Bayesian argument in regression context, and Lee et al. [2024] for the Bayesian analysis in decision-making. Another stream of literature aims to develop generalization bound or other types of convergence results. See Jeon et al. [2024], Bai et al. [2024] for such results in the regression context and Lin et al. [2023] in the decision-making context.
Pretrained transformer for RL.
The earliest effort of applying transformers to reinforcement learning lies in the area of offline RL [Janner et al., 2021, Chen et al., 2021]. By autoregressively maximizing the likelihood of trajectories in the offline dataset, this paradigm essentially converts offline RL to a supervised learning problem. When evaluating the policy, the actions are sampled according to the likelihood of PT conditioned on the so called “return-to-go”, which is commonly chosen to be the cumulative reward of “good” trajectories in the dataset. By doing so, PT retrieves information from trajectories in training data with similar return-to-go values and performs imitation learning to high-reward offline trajectories. Based on this observation, several works further examine the validity of return-to-go conditioning [Emmons et al., 2021, Ajay et al., 2022, Brandfonbrener et al., 2022] and propose alternative methods that improve this approach or extend to other settings [Laskin et al., 2022, Badrinath et al., 2024, Zheng et al., 2022, Wu et al., 2024]. For example, one of the drawbacks of the return-to-go conditioning is that it often fails in trajectory-stitching when offline data comes from sub-optimal policies.
Sequential decision-making.
We carry out experiments involving some stylized sequential decision-making models including dynamic pricing and inventory management. Dynamic pricing entails setting prices to discern the underlying revenue function, usually defined as the product of demand and price [Den Boer, 2015]. Algorithms that purely exploit based on historical data tend to be suboptimal and have been found converging to non-optimal prices with a positive probability [Lai and Robbins, 1982, Keskin and Zeevi, 2014]. Therefore, random exploration is essential for a more accurate understanding of the demand function and to avoid settling on suboptimal prices. See Harrison et al. [2012], Den Boer and Zwart [2014], Besbes and Zeevi [2015], Cohen et al. [2020] for algorithms that tackle this exploration-exploitation trade-off and achieving near-optimal regret under different settings.
For inventory management, demand estimation is different from that in dynamic pricing due to the censored demand [Huh and Rusmevichientong, 2009]. When demand exceeds inventory levels, the unmet demand is not observed, leading to lost sales. The optimal decision-making policy must balance exploration and exploitation, which involves occasionally setting inventory levels higher than usual to gather more demand data, but not so high that the additional costs of exploration become prohibitive. Regarding near-optimal policies, various works [Huh et al., 2009, Zhang et al., 2020, Agrawal and Jia, 2022] have developed algorithms that achieve an $O(\sqrt{T})$ regret bound with different assumptions about lead times.
Appendix B Problem Examples for the General Setup in Section 2
Throughout the paper, we make the following boundedness assumption.
Assumption B.1.
Assume the context space $\mathcal{X}$, the action space $\mathcal{A},$ and the observation space $\mathcal{O}$ is bounded (under the Euclidean norm) by $D$.
We provide a few examples as special cases of the framework in Section 2. We hope these examples make the actions, observations, and context in the general setup more tangible.
  • Stochastic multi-armed bandits : There is no context, $X_{t}=\text{null}$ for all $t.$ The action $a_{t}\in\mathcal{A}=\left\{1,2,\ldots,k\right\}$ denotes the index of the arm played at time $t$. The observation (also the random reward) is generated following the distribution $P_{a_{t}}$. The parameter $\gamma$ encapsulates the distributions $P_{1},\ldots,P_{k}.$
  • Linear bandits: There is no context, $X_{t}=\text{null}$ for all $t.$ The action $a_{t}\in\mathcal{A}\in\mathbb{R}^{d}$ is selected from some pre-specified domain $\mathcal{A}$. The random reward $R(X_{t},a_{t})=w^{\top}a_{t}+\epsilon_{t}$ where $\epsilon_{t}$ is some noise random variable and $w$ is a vector unknown to the agent. The observation $o_{t}=R(X_{t},a_{t})$ and the expected reward $r(X_{t},a_{t})=\mathbb{E}\left[R(X_{t},a_{t})|X_{t},a_{t}\right]=w^{\top}a_{t}$. The parameter $\gamma$ encapsulates the vector $w$ and the noise distribution.
  • Dynamic pricing: The context vector $X_{t}$ describes the market-related information at time $t$. Upon the reveal of $X_{t}$, the agent takes the action $a_{t}$ as the pricing decision. Then the agent observes the demand $d_{t}=d(X_{t},a_{t})+\epsilon_{t}$ where $\epsilon_{t}$ is some noise random variable and $d(\cdot,\cdot)$ is a demand function unknown to the agent. The random reward $R_{t}(X_{t},a_{t})=d_{t}\cdot a_{t}$ is the revenue collected from the sales at time $t$ (which equals the demand times the price), and the observation $o_{t}=d_{t}$. The expected reward $r(X_{t},a_{t})=\mathbb{E}\left[R(X_{t},a_{t})|X_{t},a_{t}\right]=d(X_{t},a_{t})\cdot a_{t}.$ The parameter $\gamma$ governs the generation of $X_{t}$’s and also contains information about the demand function $d$ and the noise distribution.
  • Newsvendor problem: As the dynamic pricing problem, the context vector $X_{t}$ describes the market-related information at time $t$. Upon its reveal, the agent takes the action $a_{t}$ which represents the number of inventory prepared for the sales at time $t$. Then the agent observes the demand $d_{t}=d(X_{t},a_{t})+\epsilon_{t}$ where $\epsilon_{t}$ is some noise random variable and $d(\cdot,\cdot)$ is a demand function unknown to the agent. The observation $o_{t}=d_{t}$, and the random reward is the negative cost $R(X_{t},a_{t})=-h\cdot(a_{t}-d_{t})^{+}-l\cdot(d_{t}-a_{t})^{+}$ where $(\cdot)^{+}$ is the positive-part function, $h$ is the left-over cost, and $l$ is the lost-sale cost. The expected reward $r(X_{t},a_{t})=\mathbb{E}\left[R(X_{t},a_{t})|X_{t},a_{t}\right]$ with the expectation taken with respect to $\epsilon_{t}$. The parameter $\gamma$ governs the generation of $X_{t}$’s and encodes the demand function $f$ and the noise distribution.
For the general problem setup in Section 2, the following proposition relates the pretraining and the regret through the lens of loss function.
Proposition B.2 (Surrogate property).
We say the loss function $l(\cdot,\cdot)$ satisfies the surrogate property if there exists a constant $C>0$ such that
$\mathrm{Regret}(\texttt{TF}_{\theta};\gamma)\leq C\cdot L\left(\texttt{TF}_{\theta};\gamma\right)$
holds for any $\texttt{TF}_{\theta}$ and $\gamma$. We have
  • (a)
    The cross-entropy loss satisfies the surrogate property for the multi-armed bandits problem.
  • (b)
    The squared loss satisfies the surrogate property for the dynamic pricing problem.
  • (c)
    The absolute loss satisfies the surrogate property for the linear bandits and the newsvendor problem.
Proposition B.2 gives a first validity of the supervised pretraining approach. The surrogate property is a property of the loss function and it is not affected by the distribution or the decision function $\texttt{TF}_{\theta}.$ Specifically, it upper bounds the regret with the action prediction loss. It is also natural in that a closer prediction of the optimal action gives a smaller regret. One implication is that the result guides the choice of the loss function for different underlying problems.
B.1 Proof of Proposition B.2
To prove Proposition B.2, without loss of generality, we assume the following mild conditions. First, for all these problems, the noise random variable $\epsilon_{t}$ has mean 00 and is independent of the action $a_{t}$. Second, for the linear bandits problem, the $L_{\infty}$ norm of the unknown $w$ is bounded by $D$. Third, for the dynamic pricing problem and the newsvendor problem, we assume the demand function is differentiable with respect to the action coordinate, and the first and second-order derivatives are $D$-Lipschitz.
Proof.
Now we prove Proposition B.2 part by part.
(a). For the multi-arm bandits problem, let the optimal arm is the first one. For $1\leq i\leq k$, let $\mu_{i}$ be the mean of the reward distribution for arm $i$, and denote by $\Delta_{i}=\mu_{1}-\mu_{i}$ the action gap for arm $i$. Let $\Delta_{\max}=\max\left\{\Delta_{1},\cdots,\Delta_{k}\right\}$ be the maximum action gap.
Since $\texttt{TF}_{\theta}(H_{t})$ can be viewed as a distribution over $\left\{1,\cdots,k\right\}$, we now rewrite $P(\texttt{TF}_{\theta}(H_{t})=i)$ as $p^{t}_{i}$ for notation simplicity. We claim that $\text{Regret}(\texttt{TF}_{\theta};\gamma)\leq\Delta_{\max}\cdot L(\texttt{TF}_{\theta};\gamma)$, where $L(\cdot,\cdot)$ is the cross entropy loss. To see this, observe that condition on $H_{t}$,
$\mathbb{E}\left[r\left(X_{t},a_{t}^{*}\right)-r\left(X_{t},a_{t}\right)|H_{t}\right]$ $=\mu_{1}-\sum_{i=1}^{k}p^{t}_{i}\mu_{i}=\sum_{i=2}^{k}p^{t}_{i}\Delta_{i}\leq\Delta_{\max}(1-p^{t}_{1})\leq-\Delta_{\max}\cdot\log(p^{t}_{1}),$
then take the expectation of both sides and take the sum from $t=1$ to $t=T$ of both sides completes the proof.
(b). For the dynamic pricing problem, we will show that $\text{Regret}(\texttt{TF}_{\theta};\gamma)\leq D\cdot L(\texttt{TF}_{\theta};\gamma)$, where $L(\cdot,\cdot)$ is the squared $L_{2}$ loss. From Taylor’s expansion, we have for any $a_{t}\in\mathcal{A}$,
$r(X_{t},a_{t})-r(X_{t},a^{*}_{t})\geq-D||a_{t}-a_{t}^{*}||_{2}^{2},$
which implies that $\mathbb{E}\left[r(X_{t},a_{t}^{*})-r(X_{t},a_{t})\right]\leq D||a_{t}-a_{t}^{*}||_{2}^{2}.$ Then taking the sum from $t=1$ to $t=T$ gives the results.
[…]
$\mathbb{E}\left[r\left(X_{t},a_{t}^{*}\right)-r\left(X_{t},a_{t}\right)\right]=\mathbb{E}\left[w^{T}a^{*}-w^{T}a_{t}\right]\leq\mathbb{E}\left[||w||_{\infty}||a^{*}-a_{t}||_{1}\right]\leq D\cdot\mathbb{E}\left[||a_{t}^{*}-f_{\theta}(H_{t})||_{1}\right].$
[…]
$r(X_{t},a_{t}^{*})-r(X_{t},a_{t})$ $\leq\max\left\{h,l\right\}|a_{t}^{*}+d(X_{t},a_{t}^{*})-a_{t}-d(X_{t},a_{t})|$
$=\max\left\{h,l\right\}\left(\mathbf{1}+\nabla_{a}d(X_{t},\xi)\right)^{T}(a_{t}^{*}-a_{t})$
$\leq(D+1)\max\left\{h,l\right\}||a_{t}^{*}-a_{t}||_{1}.$
Then taking the expectation and taking the sum from $t=1$ to $t=T$ of both sides gives the result. ∎
Appendix C More Experiments
C.1 More experiments on training dynamics
In addition to Figure 1 (a), we further evaluate the effectiveness of Algorithm 1 across various tasks and architectures, as illustrated in Figure 4.
(a) Linear bandits
(b) Newsvendor
(c) Multi-armed bandits (DPT)
Figure 4: Effectiveness of Algorithm 1. The values are averaged out-of-sample regrets based on $128$ runs, where the shaded area indicates the standard deviation.
Figure 4 presents comparisons of the transformer’s testing regret (see definition and implementation in Appendix E.1) with and without self-generated data (i.e., using Algorithm 1 or not) during the training, colored by blue and orange respectively. The comparisons are conducted on (a) linear bandits and (b) newsvendor tasks. Additionally, (c) includes a similar comparison conducted in a model with different architecture, the Decision Pre-trained Transformer (DPT) Lee et al. [2024], on the multi-armed bandits task. The results consistently demonstrate the same pattern observed in Figure 1 (a): incorporating transformer-generated data significantly reduces the testing regret loss at each iteration $m>50$ compared to not using it. This shows the effectiveness of Algorithm 1 (i) across different tasks and (ii) for various model architectures.
Setup. In Figure 4 (a) and (b), we consider two tasks: a linear bandits task (with $2$-dimensional action space) and a newsvendor task (with $4$-dimensional contexts). Both tasks have infinite support of the prior environment distribution $\mathcal{P}_{\gamma}$ (i.e., infinite possible environments in both the pretraining and testing). The details of the tasks can be found in Appendix E.2. For each task, we independently train two transformer models by Algorithm 1. For both models, the training parameters are identical except for the $M_{0}$: the first curve (blue, thick line) uses $M_{0}=50$, while the second curve (orange, dashed line) uses $M_{0}=130$, meaning no transformer-generated data is utilized during training. All other parameters follow the configuration detailed in Appendix E.1. The figure shows the mean testing regret at each training iteration across 128 randomly sampled environments from $\mathcal{P}_{\gamma}$, with the shaded areas representing the standard deviation. We conduct the same experiment on the multi-armed bandits task (with infinite possible environments and $20$ arms) for a different architecture for sequential decision making, Decision Pretrained Transformer (DPT) [Lee et al., 2024], in Figure 4 (c). We refer to the original paper for details on the DPT’s architecture.
C.2 More experiments on matchings of $\texttt{TF}_{\hat{\theta}}$ and $\texttt{Alg}^{*}$
This section provides more results regarding the matching of $\texttt{TF}_{\hat{\theta}}$ and $\texttt{Alg}^{*}$ across different tasks: Figure 5 provides more examples comparing decisions made by $\texttt{TF}_{\hat{\theta}}$ and $\texttt{Alg}^{*}$ at sample path levels, while Figure 6 compares them at population levels.
(a) Dynamic pricing, $100$ environments, linear demand type
(b) Dynamic pricing, $16$ environments, linear and square demand types
(c) Newsvendor, $100$ environments, linear demand type
(d) Newsvendor, $16$ environments, linear and square demand types
Figure 5: Examples to compare the actions from the transformer $\texttt{TF}_{\hat{\theta}}$ and the optimal decision function $\texttt{Alg}^{*}$.
(a) Multi-armed bandits, 4 environments
(b) Dynamic pricing, 4 environments
(c) Newsvendor, $4$ environments
(d) Linear bandits, 4 environments
(e) Multi-armed bandits, 100 environments
(f) Dynamic pricing, 100 environments
(g) Newsvendor, 100 environments
(h) Linear bandits, 100 environments
Figure 6: $\texttt{TF}_{\hat{\theta}}(h_{t})-\texttt{Alg}^{*}(h_{t})$ across different tasks with various numbers of possible environments. More possible environments lead to a harder decision-making problem.
We plot the actions generated from $\texttt{TF}_{\hat{\theta}}$ and $\texttt{Alg}^{*}$ on the same environment with the same contexts in Figure 5 across dynamic pricing and newsvendor tasks. Each subfigure has a different number of possible environments or possible demand types (which are shared in both the pretraining and testing phases). The population-level differences $\{\texttt{TF}_{\hat{\theta}}(h_{t})-\texttt{Alg}^{*}(h_{t})\}_{t=1}^{T}$ are presented in Figure 6. We observe that across different tasks: (1) $\texttt{TF}_{\hat{\theta}}$ nearly matches $\texttt{Alg}^{*}$, but (2) the matchings are not perfect. The actions from $\texttt{TF}_{\hat{\theta}}$ and $\texttt{Alg}^{*}$ are not exactly the same, and their differences increase as the underlying problems become more complex, as seen by comparing the first and second rows in Figure 6. This latter observation numerically supports the conditions in Proposition 4.4.
Setup. For Figure 5, each subfigure is based on a sampled environment with a sampled sequence of contexts $\{X_{t}\}_{t=1}^{30}$ from the corresponding task. For tasks in (a),(c), the data generation process follows the description detailed in Appendix E.2, and the architecture of $\texttt{TF}_{\hat{\theta}}$ and the definition of $\texttt{Alg}^{*}$ can be found in Appendix E.1 and Appendix E.3 respectively. For tasks in (b), (d), there are 16 environments included in the support of $\mathcal{P}_{\gamma}$, where 8 environments have demand functions of the linear type and the other 8 have demand functions of the square type (see Appendix E.2 for the definitions of these two types). Both the pretraining and testing samples are drawn from $\mathcal{P}_{\gamma}$, i.e., from these 16 environments. The remaining setups for these tasks follow the methods provided in Appendix E.2.
For each task shown in Figure 6, we generate $100$ sequences of $\{\texttt{TF}_{\hat{\theta}}(h_{t})-\texttt{Alg}^{*}(h_{t})\}_{t=1}^{100}$, i.e., each subfigure is based on $10000$ samples of $\texttt{TF}_{\hat{\theta}}(h_{t})-\texttt{Alg}^{*}(h_{t})$. For dynamic pricing and newsvendor, actions are scalars and thus $\texttt{TF}_{\hat{\theta}}(h_{t})-\texttt{Alg}^{*}(h_{t})$ is also a scalar; For multi-armed bandits, the action space is discrete and we use the reward difference associated with the chosen actions (by $\texttt{TF}_{\hat{\theta}}(h_{t})$ or $\texttt{Alg}^{*}(h_{t})$) as the value of $\texttt{TF}_{\hat{\theta}}(h_{t})-\texttt{Alg}^{*}(h_{t})$. We use histograms to summarize the samples from them. For linear bandits, the actions are $2$-dimensional and thus we use the scatters to show $\texttt{TF}_{\hat{\theta}}(h_{t})-\texttt{Alg}^{*}(h_{t})$ in Figure 6 (d) or (h).
C.3 Comparison with oracle posterior algorithms
In this subsection, we study the behavior of $\texttt{TF}_{\hat{\theta}}$ by comparing regret performances on $\texttt{TF}_{\hat{\theta}}$ and the algorithms utilizing the posterior distribution $\mathcal{P}(\gamma|H)$ of environments. We name the later algorithms as the oracle posterior algorithms as they are assumed to know the posterior distribution $\mathcal{P}(\gamma|H)$ (or equally, the prior distribution $\mathcal{P}(\gamma)$).
We first present a general framework for these oracle posterior algorithms in Algorithm 2: posterior averaging, sampling, and median. For simplicity of notation, we omit the subscript $t$ and only consider a finite space $\Gamma$ of environments, while the algorithms can be easily extended to the infinite case. For the posterior median algorithm, we assume the action space $\mathcal{A}\subseteq\mathbb{R}$.
Algorithm 2 Posterior averaging/sampling/median at time $t$
[…]
2:    %Posterior averaging
3:if Alg$=$posterior averaging then
4:     $a=\sum_{\gamma\in\Gamma}\mathcal{P}(\gamma|H)\cdot a^{*}_{\gamma}$
5:    %Posterior sampling
6:else if Alg$=$posterior sampling then
7:     $a=a^{*}_{\tilde{\gamma}}$, where $\tilde{\gamma}\sim\mathcal{P}(\cdot|H)$
8:    %Posterior median
9:else if Alg$=$posterior median then
10:     Sort and index the environments as $\{\gamma_{i}\}_{i=1}^{|\Gamma|}$ such that the corresponding optimal actions are in ascending order: $a^{*}_{\gamma_{1}}\leq a^{*}_{\gamma_{2}}\leq\ldots\leq a^{*}_{\gamma_{|\Gamma|}}$.
11:     Choose $a=\min\left\{a^{*}_{\gamma_{i}}\big{|}\sum_{i^{\prime}=1}^{i}\mathcal{P}(\gamma_{i^{\prime}}|H)\geq 0.5\right\}$
12:end if
13:Return: $a$
(a) Regret for multi-armed bandits
(b) Action suboptimality for multi-armed bandits
(c) Regret for linear bandits
(d) Action suboptimality for linear bandits
Figure 7: Performances of $\texttt{TF}_{\hat{\theta}}$ and oracle posterior algorithms. The numbers in the legend bar are the final regret at $t=100$ and the shaded areas indicate the $90\%$ (empirical) confidence intervals.
Figure 7 presents the testing regret and action suboptimality for $\texttt{TF}_{\hat{\theta}}$ compared to oracle posterior algorithms. Generally, all algorithms exhibit good regret performance. However, we observe that $\texttt{TF}_{\hat{\theta}}$ is (i) superior to posterior sampling in multi-armed bandits, as shown in (a), and (ii) outperforms posterior averaging in linear bandits, as shown in (c).
Specifically, (b) shows that posterior sampling has higher action suboptimality during the initial time steps compared to $\texttt{TF}_{\hat{\theta}}$. This suggests that the exploration inherent in the sampling step of posterior sampling can introduce additional regret, which might be unnecessary for simpler problems. Conversely, (d) indicates that the decisions from posterior averaging do not converge to the optimal action $a^{*}_{t}$ as quickly as those from $\texttt{TF}_{\hat{\theta}}$ and posterior sampling. This suggests that posterior averaging may be too greedy, thus failing to sufficiently explore the environment. This second observation also numerically supports Proposition 4.3.
These observations illustrate that $\texttt{TF}_{\hat{\theta}}$ can outperform oracle posterior algorithms in certain scenarios. This further demonstrates that the advantage of $\texttt{TF}_{\hat{\theta}}$ over benchmark algorithms is not solely due to its use of prior knowledge about the task. Rather, $\texttt{TF}_{\hat{\theta}}$ discovers a new decision rule that achieves better short-term regret than oracle posterior algorithms. It can be more greedy than posterior sampling while exploring more than posterior averaging.
Setup. We consider a multi-armed bandits task (with $20$ arms) and linear bandits task (with $2$-dimensional actions) with $4$ environments. The results are based on $100$ runs. The details of tasks with their specific oracle posterior algorithms are provided in Appendix E.2 and Appendix E.3.
C.4 $\texttt{TF}_{\hat{\theta}}$ as a solution to model misspecification
Most sequential decision-making algorithms, including all the benchmark algorithms in our experiments, rely on structural or model assumptions about the underlying tasks. For instance, demand functions in pricing and newsvendor problems are typically assumed to be linear in context. Applying these algorithms in misspecified environments, where these assumptions do not hold, can lead to degraded performance.
In contrast, $\texttt{TF}_{\hat{\theta}}$ offers a potential solution to model misspecifications. By generating pretraining samples from all possible types of environments (e.g., newsvendor models with both linear and non-linear demand functions), $\texttt{TF}_{\hat{\theta}}$ leverages its large capacity to make near-optimal decisions across different types. Figure 8 demonstrates the performance of $\texttt{TF}_{\hat{\theta}}$ trained on tasks with two types of demand functions: linear and square (definitions provided in Appendix E.2). It compares $\texttt{TF}_{\hat{\theta}}$ with benchmark algorithms designed solely for linear demand cases. The performance is tested in environments where demand functions (i) can be either type with equal probability (first column); (ii) are strictly square type (second column); and (iii) are strictly linear type (third column). The benchmark algorithms face model misspecification issues in the first two cases.
For both pricing and newsvendor tasks, $\texttt{TF}_{\hat{\theta}}$ consistently outperforms all benchmarks across the three scenarios, especially in the first two cases. This result highlights the potential of $\texttt{TF}_{\hat{\theta}}$ to effectively handle model misspecifications.
(a) Pricing, two demands
(b) Pricing, square demand
(c) Pricing, linear demand
(d) Newsvendor, two demands
(e) Newsvendor, square demand
(f) Newsvendor, linear demand
Figure 8: The average out-of-sample regret on dynamic pricing (first row) and newsvendor (second row) with two possible demand types: linear and square. The numbers in the legend bar are the final regret at $t=100$ and the shaded areas indicate the $90\%$ (empirical) confidence intervals. The details of benchmarks can be found in Appendix E.3.
Setup. For both the pricing and newsvendor tasks, during the pretraining phase of $\texttt{TF}_{\hat{\theta}}$, the pretraining data are generated from two types of demand functions: linear and square, with each type having a half probability. The support of $\mathcal{P}_{\gamma}$ is infinite, meaning the parameters in the demand functions are not restricted. The context dimension is 6 for the pricing task and 4 for the newsvendor task. The generating details and definitions of these two types of demands are provided in Appendix E.2. Additionally, when generating data using $\texttt{TF}_{\theta_{m}}$ in Algorithm 1, the sampled environment also has an equal probability of having a linear or square demand type.
Each subfigure in Figure 8 is based on 100 runs. We consider three different cases during testing: (i) the environment can have either the square or linear demand type (each with half probability); (ii) the environment can only have the linear demand type; and (iii) the environment can only have the square demand type. In the first case, the testing environment is sampled in the same way as the pretraining environment, while in the other two cases, we only include the sampled environments with the specified demand type until the total testing samples reach 100.
C.5 Additional results on $\texttt{TF}_{\hat{\theta}}$ performances
This subsection presents further results on the performance of $\texttt{TF}_{\hat{\theta}}$. Specifically, we compare $\texttt{TF}_{\hat{\theta}}$ with benchmark algorithms on (i) simple tasks, which include only 4 possible environments in both pretraining and testing samples (Figure 9), and (ii) more complex tasks, which involve either 100 possible environments (Figure 10 (a), (b)) or 16 environments with two possible types of demand functions (Figure 10 (c), (d)).
Figures 9 and 10 illustrate the consistently and significantly superior performance of $\texttt{TF}_{\hat{\theta}}$ and $\texttt{Alg}^{*}$ across all tasks compared to the benchmark algorithms. These results underscore the advantage of leveraging prior knowledge about the tested environments.
(a) Multi-armed bandits
(b) Linear bandits
(c) Dynamic pricing
(d) Newsvendor
Figure 9: The average out-of-sample regret on tasks with simpler environments, where each task only has $4$ possible environments. The numbers in the legend bar are the final regret at $t=100$ and the shaded areas indicate the $90\%$ (empirical) confidence intervals. The details of benchmarks can be found in Appendix E.3.
(a) Multi-armed bandits, 100 environments, linear demand
(b) Linear bandits, 100 environments, linear demand
(c) Dynamic pricing, 16 environments, 2 demand types
(d) Newsvendor, 16 environments, 2 demand types
Figure 10: The average out-of-sample regret on harder tasks: with 100 environments in (a) multi-armed bandits and (b) linear bandits; or with 16 environments, 2 demand types in (c) dynamic pricing and (d) newsvendor. The numbers in the legend bar are the final regret at $t=100$ and the shaded areas indicate the $90\%$ (empirical) confidence intervals. The details of benchmarks can be found in Appendix E.3
Setup. All figures are based on 100 runs. We consider four tasks with 4 environments (i.e., the support of $\mathcal{P}_{\gamma}$ contains only 4 environments) in Figure 9 and two tasks with 100 environments in Figure 10 (a), (b). The setup for these tasks follows the generation methods provided in Appendix E.2.
For the tasks in Figure 10 (c), (d), there are 16 environments included in the support of $\mathcal{P}_{\gamma}$, where 8 environments have demand functions of the linear type and the other 8 have demand functions of the square type (see Appendix E.2 for the definitions of these two types). Both the pretraining and testing samples are drawn from $\mathcal{P}_{\gamma}$, i.e., from these 16 environments. The remaining setups for these tasks follow the methods provided in Appendix E.2.
Appendix D Proofs
D.1 Discussions on Claim 3.1
Consider the following setup where a feature-target pair $(X,Y)$ is generated from some distribution $\mathcal{P}_{X,Y}(\theta)$. Then we consider the minimization of the following loss
$\min_{\theta}\mathbb{E}_{\mathcal{P}_{X,Y}(\theta)}[l(f_{\theta}(X),Y)]$
for some loss function $l:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}.$
Under this formulation, both the data generation distribution $\mathcal{P}_{X,Y}(\theta)$ and the prediction function $f_{\theta}$ are parameterized by $\theta$. This describes the performative prediction problem [Perdomo et al., 2020].
In addition, we note that in our pretrained transformer setting, the joint distribution can be factorized in
$\mathcal{P}_{X,Y}(\theta)=\mathcal{P}_{X}(\theta)\cdot\mathcal{P}_{Y|X}.$
That is, the parameter $\theta$ only induces a covariate shift by affecting the marginal distribution of $X$, but the conditional distribution $\mathcal{P}_{Y|X}$ remains the same for all the $\theta$. This exactly matches our setting of pretrained transformer; to see this, different decision functions, $f$ or $\texttt{TF}_{\theta_{m}}$, only differ in terms of the generation of the actions $a_{\tau}$’s in $H_{t}$ (for $\tau=1,...,t-1$) which is the features $X$ in this formulation but the optimal action $a_{t}^{*}$ will only be affected by $X_{t}$ in $H_{t}$.
Also, under this factorization, we define the Bayes-optimal estimator as
$f^{*}(X)=\min_{y}\mathbb{E}_{\mathcal{P}_{Y|X}}[l(y,Y)|X].$
Then it is easy to note that when $f^{*}=f_{\theta^{*}}$ for some $\theta^{*}$, and have
$\theta^{*}\in\operatorname*{arg\,min}_{\theta}\mathbb{E}_{\mathcal{P}_{X,Y}(\theta^{\prime})}[l(f_{\theta}(X),Y)]$
for any $\theta^{\prime}.$
In this light, the oscillating behavior of the optimization algorithms in Perdomo et al. [2020] will not happen. Because for all the data generation distribution $\mathcal{P}_{X,Y}(\theta^{\prime})$, they all point to one optimal $\theta^{*}$.
Back to the context of the pretrained transformer, such a nice property is contingent on two factors: (i) the transformer function class is rich enough to cover $f^{*}$; (ii) there are infinitely many samples/we can use the expected loss. Also, the above argument is connected to the argument of Proposition 4.1.
D.2 Proof of Proposition 3.2
As noted in the proposition, let $b$ denote the bound of the loss function, and let the loss function $l\left(\texttt{TF}_{{\theta}}(H),a\right)$ is $D$-Lipschitz with respect to $H$ and $a$. We remark that the constant $D$ could be very large for the transformer model; the bound derived here serves more to illustrate the relationship between the generalization error and various problem parameters, but not really as an empirical bound to predict the test error of the underlying transformer.
Step 1. We aim to show with probability at least $1-h$,
$\kappa L_{f}(\texttt{TF}_{\hat{\theta}})+(1-\kappa)L_{\texttt{TF}_{\tilde{\theta}}}(\texttt{TF}_{\hat{\theta}})\leq\frac{1}{nT}\mathop{\sum}\limits_{i=1}^{n}\mathop{\sum}\limits_{t=1}^{T}l\left(\texttt{TF}_{\hat{\theta}}\left(H_{t}^{(i)}\right),a_{t}^{(i)*}\right)+\sqrt{\frac{\text{Complex}(\{\texttt{TF}_{\theta}:\theta\in\Theta\})}{nT}}+\sqrt{\frac{2b^{2}}{nT}\log\left(\frac{4}{h}\right)},$
where $\text{Complex}(\{\texttt{TF}_{\theta}:\theta\in\Theta\})$ is some inherent constant describing the complexity of $\{\texttt{TF}_{\theta}:\theta\in\Theta\}$. Note that when samples follow the generation process (actions are generated by $f$)
$\mathcal{P}_{\gamma}\rightarrow\gamma_{i}\rightarrow\mathcal{P}_{\gamma_{i},f}\rightarrow\left\{\left(H_{1}^{(i)},a_{1}^{(i)*}\right),\ldots,\left(H_{T}^{(i)},a_{T}^{(i)*}\right)\right\},$
and
$L_{f}(\texttt{TF}_{\theta})=\mathbb{E}_{\gamma\sim\mathcal{P}_{\gamma}}\left[L_{f}(\texttt{TF}_{\theta};\gamma)\right]=\mathbb{E}_{\gamma\sim\mathcal{P}_{\gamma}}\left[\mathbb{E}_{(H_{t},a_{t}^{*})\sim\mathcal{P}_{\gamma,f}}\left[\sum_{t=1}^{T}l\left(\texttt{TF}_{\theta}\left(H_{t}\right),a_{t}^{*}\right)\right]\right].$
[…]
$\mathop{\sup}\limits_{\theta\in\Theta}\Bigg{|}\frac{1}{nT}\mathop{\sum}\limits_{i=1}^{n}\mathop{\sum}\limits_{t=1}^{T}l\left(\texttt{TF}_{\theta}\left(H_{t}^{(i)}\right),a_{t}^{(i)*}\right)-L_{f}(\texttt{TF}_{\theta})\Bigg{|}$
$\leq$ $\mathbb{E}_{(H_{t},a_{t}^{*})\sim\mathcal{P}_{\gamma,f}}\left[\mathop{\sup}\limits_{\theta\in\Theta}\Bigg{|}\frac{1}{nT}\mathop{\sum}\limits_{i=1}^{n}\mathop{\sum}\limits_{t=1}^{T}l\left(\texttt{TF}_{\theta}\left(H_{t}^{(i)}\right),a_{t}^{(i)*}\right)-\mathbb{E}_{(H_{t},a_{t}^{*})\sim\mathcal{P}_{\gamma,f}}\left[\mathop{\sum}\limits_{t=1}^{T}l\left(\texttt{TF}_{\theta}\left(H_{t}\right),a_{t}^{*}\right)\right]\Bigg{|}\right]+\sqrt{\frac{2b^{2}}{nT}\log\left(\frac{2}{h}\right)}$
$\leq$ $2\mathbb{E}_{(H_{t},a_{t}^{*})\sim\mathcal{P}_{\gamma,f},w_{it}}\left[\mathop{\sup}\limits_{\theta\in\Theta}\Bigg{|}\frac{1}{nT}\mathop{\sum}\limits_{i=1}^{n}\mathop{\sum}\limits_{t=1}^{T}w_{it}l\left(\texttt{TF}_{\theta}\left(H_{t}^{(i)}\right),a_{t}^{(i)*}\right)\Bigg{|}\right]+\sqrt{\frac{2b^{2}}{nT}\log\left(\frac{2}{h}\right)}$
$\leq$ $\sqrt{\frac{\text{Complex}(\{\texttt{TF}_{\theta}:\theta\in\Theta\})}{nT}}+\sqrt{\frac{2b^{2}}{nT}\log\left(\frac{2}{h}\right)},$ (10)
where the second inequality uses the symmetric tricks (see Theorem 3.3 in [Mohri et al., 2018] for more details), and $w_{it}$, so-called Rademacher variables are independent uniform random variables taking values in $\left\{-1,1\right\}$.
Similarly, when samples follow the generation process (actions are generated by $\text{TF}_{\bar{\theta}}$)
$\mathcal{P}_{\gamma}\rightarrow\gamma_{i}\rightarrow\mathcal{P}_{\gamma_{i},\texttt{TF}_{\tilde{\theta}}}\rightarrow\left\{\left(H_{1}^{(i)},a_{1}^{(i)*}\right),\ldots,\left(H_{T}^{(i)},a_{T}^{(i)*}\right)\right\},$
by using the same tricks, we can have that with probability at least $1-h$,
$\mathop{\sup}\limits_{\theta\in\Theta}\Bigg{|}\frac{1}{nT}\mathop{\sum}\limits_{i=1}^{n}\mathop{\sum}\limits_{t=1}^{T}l\left(\texttt{TF}_{\theta}\left(H_{t}^{(i)}\right),a_{t}^{(i)*}\right)-L_{\texttt{TF}_{\tilde{\theta}}}(\texttt{TF}_{\theta})\Bigg{|}\leq\sqrt{\frac{\text{Complex}(\{\texttt{TF}_{\theta}:\theta\in\Theta\})}{nT}}+\sqrt{\frac{2b^{2}}{nT}\log\left(\frac{2}{h}\right)}.$ (11)
[…]
$\kappa L_{f}(\texttt{TF}_{\hat{\theta}})+(1-\kappa)L_{\texttt{TF}_{\tilde{\theta}}}(\texttt{TF}_{\hat{\theta}})\leq\frac{1}{nT}\mathop{\sum}\limits_{i=1}^{n}\mathop{\sum}\limits_{t=1}^{T}l\left(\texttt{TF}_{\hat{\theta}}\left(H_{t}^{(i)}\right),a_{t}^{(i)*}\right)+\sqrt{\frac{\text{Complex}(\{\texttt{TF}_{\theta}:\theta\in\Theta\})}{nT}}+\sqrt{\frac{2b^{2}}{nT}\log\left(\frac{4}{h}\right)}.$
Step 2. Now we show
$L(\texttt{TF}_{\hat{\theta}})\leq L_{f}(\texttt{TF}_{\hat{\theta}})+TD\cdot\mathbb{E}_{\gamma\sim\mathcal{P}_{\gamma}}\left[W_{1}\left(\mathcal{P}_{\gamma,f},\mathcal{P}_{\gamma,\texttt{TF}_{\hat{\theta}}}\right)\right],\ L(\texttt{TF}_{\hat{\theta}})\leq L_{f}(\texttt{TF}_{\hat{\theta}})+TD\cdot\mathbb{E}_{\gamma\sim\mathcal{P}_{\gamma}}\left[W_{1}\left(\mathcal{P}_{\gamma,\texttt{TF}_{\tilde{\theta}}},\mathcal{P}_{\gamma,\texttt{TF}_{\hat{\theta}}}\right)\right],$
where $W_{1}(\cdot,\cdot)$ is the Wasserstein-1 distance.
By Kantorovich-Rubinstein’s theorem, for distributions $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, we have
[…]
Since $l\left(\texttt{TF}_{{\theta}}(H),a\right)$ is $D$-Lipschitz with respect to $H$ and $a$ for all $\theta\in\Theta$, we have
$L\left(\texttt{TF}_{\hat{\theta}};\gamma\right)-L_{f}\left(\texttt{TF}_{\hat{\theta}};\gamma\right)$ $=\mathbb{E}_{(H_{t},a_{t}^{*})\sim\mathcal{P}_{\gamma,\texttt{TF}_{\hat{\theta}}}}\left[\sum_{t=1}^{T}l\left(\texttt{TF}_{\hat{\theta}}\left(H_{t}\right),a_{t}^{*}\right)\right]-\mathbb{E}_{(H_{t},a_{t}^{*})\sim\mathcal{P}_{\gamma,f}}\left[\sum_{t=1}^{T}l\left(\texttt{TF}_{\hat{\theta}}\left(H_{t}\right),a_{t}^{*}\right)\right]$
$\leq TD\cdot W_{1}\left(\mathcal{P}_{\gamma,f},\mathcal{P}_{\gamma,\texttt{TF}_{\hat{\theta}}}\right),$
which implies that $L(\texttt{TF}_{\hat{\theta}})\leq L_{f}(\texttt{TF}_{\hat{\theta}})+TD\cdot\mathbb{E}_{\gamma\sim\mathcal{P}_{\gamma}}\left[W_{1}\left(\mathcal{P}_{\gamma,f},\mathcal{P}_{\gamma,\texttt{TF}_{\hat{\theta}}}\right)\right]$. Similarly, we have $L(\texttt{TF}_{\hat{\theta}})\leq L_{f}(\texttt{TF}_{\hat{\theta}})+TD\cdot\mathbb{E}_{\gamma\sim\mathcal{P}_{\gamma}}\left[W_{1}\left(\mathcal{P}_{\gamma,\texttt{TF}_{\tilde{\theta}}},\mathcal{P}_{\gamma,\texttt{TF}_{\hat{\theta}}}\right)\right]$.
Therefore, by combining step 1 and step 2, we can conclude that with probability at least $1-h$,
$L(\texttt{TF}_{\hat{\theta}})\leq$ $\hat{L}(\texttt{TF}_{\hat{\theta}})+\sqrt{\frac{\text{Complex}(\{\texttt{TF}_{\theta}:\theta\in\Theta\})}{nT}}$
$+$ $\kappa TD\cdot\mathbb{E}_{\gamma\sim\mathcal{P}_{\gamma}}\left[W_{1}\left(\mathcal{P}_{\gamma,f},\mathcal{P}_{\gamma,\texttt{TF}_{\hat{\theta}}}\right)\right]+(1-\kappa)TD\cdot\mathbb{E}_{\gamma\sim\mathcal{P}_{\gamma}}\left[W_{1}\left(\mathcal{P}_{\gamma,\texttt{TF}_{\tilde{\theta}}},\mathcal{P}_{\gamma,\texttt{TF}_{\hat{\theta}}}\right)\right]+\sqrt{\frac{2b^{2}}{nT}\log\left(\frac{4}{h}\right)},$
which completes the proof.
D.3 Proof of Proposition 4.1
Proof.
The proof can be done with the definition of $\texttt{Alg}^{*}.$ Also, we refer more discussion to Section D.1. ∎
D.4 Examples for the optimal decision function $\texttt{Alg}^{*}$ in Section 4
Here we provide some additional examples in each problem context to generate more intuitions of the optimal decision function $\texttt{Alg}^{*}$ in Section 4.
(1). Multi-armed bandits under the cross-entropy loss: For convenience, we rewrite the best arm that corresponds to environment $\gamma$, i.e., $a^{*}(H_{t},\gamma)=\mathop{\operatorname*{arg\,max}}\limits_{i}\ \mathbb{E}\left[o_{t}|\gamma_{i}\right]$ as $i_{\gamma}$. Then,
$\texttt{Alg}^{*}(H_{t})$ $=\operatorname*{arg\,min}_{\sum_{i=1}^{k}p_{i}^{t}=1}\int_{\gamma}-\log\left(p_{i_{\gamma}}^{t}\right)\mathbb{P}(H_{t}|\gamma)\mathrm{d}\mathcal{P}_{\gamma}=\operatorname*{arg\,min}_{\sum_{i=1}^{k}p_{i}^{t}=1}\ \sum_{j=1}^{k}\left(-\log\left(p_{j}^{t}\right)\right)\int_{\left\{\gamma:\ i_{\gamma}=j\right\}}\mathbb{P}(H_{t}|\gamma)\mathrm{d}\mathcal{P}_{\gamma}.$
By nothing $\int_{\left\{\gamma:\ i_{\gamma}=j\right\}}\mathbb{P}(H_{t}|\gamma)\mathrm{d}\mathcal{P}_{\gamma}$ is the probability such that action $j$ is the optimal action conditional on $H_{t}$ (i.e., the posterior distribution of the optimal action), we are indeed minimizing the cross-entropy of the posterior distribution of the optimal action relative to the decision variables $p_{j}$. Thus, the optimal solution $p^{*}_{j}=\int_{\left\{\gamma:\ i_{\gamma}=j\right\}}\mathbb{P}(H_{t}|\gamma)\mathrm{d}\mathcal{P}_{\gamma}$ for each $j$ and $\texttt{Alg}^{*}(H_{t})$ behaves as the posterior sampling.
(2). Linear bandits under the squared $L_{2}$ loss: We rewrite the best arm that corresponds to environment $\gamma$, i.e., $a^{*}(H_{t},\gamma)=\mathop{\operatorname*{arg\,max}}\limits_{a\in\mathcal{A}}\ w_{\gamma}^{T}a$ as $a_{(\gamma)}$. Then
$\texttt{Alg}^{*}(H_{t})$ $=\operatorname*{arg\,min}_{a\in\mathcal{A}}\int_{\gamma}||a-a^{*}(H_{t},\gamma)||^{2}\mathbb{P}(H_{t}|\gamma)\mathrm{d}\mathcal{P_{\gamma}}=\operatorname*{arg\,min}_{a\in\mathcal{A}}\int_{\gamma}||a-a_{(\gamma)}||^{2}\mathbb{P}(H_{t}|\gamma)\mathrm{d}\mathcal{P_{\gamma}}$
$=\frac{\int_{\gamma}a_{(\gamma)}\mathbb{P}(H_{t}|\gamma)\mathrm{d}\mathcal{P_{\gamma}}}{\int_{\gamma}\mathbb{P}(H_{t}|\gamma)\mathrm{d}\mathcal{P_{\gamma}}}=\int_{\gamma}a_{(\gamma)}\ \mathrm{d}\mathcal{P}\left(\gamma|H_{t}\right),$
where $\mathcal{P}\left(\gamma|H_{t}\right)$ is the posterior distribution of $\gamma$. Notice that the second last equation holds by taking the derivative of $a$ and applying the first-order condition. Hence $\texttt{Alg}^{*}(H)$ is the posterior averaging.
(3). Newsvendor problem under the $L_{1}$ loss. Let
$a^{*}(H_{t},\gamma)=\mathop{\operatorname*{arg\,max}}\limits_{a\in\mathbb{R}}\ r(X_{t},a)=\mathop{\operatorname*{arg\,max}}\limits_{a\in\mathbb{R}}\ \mathbb{E}\left[-h\cdot(a_{t}-o_{t})^{+}-l\cdot(o_{t}-a_{t})^{+}|X_{t},a_{t}\right]\triangleq a_{(\gamma)}.$
Then by zero-subgradient condition, we can conclude that $\texttt{Alg}^{*}(H_{t})=\mathop{\operatorname*{arg\,min}}\limits_{a\in\mathbb{R}}\int_{\gamma}|a-a_{(\gamma)}|\ \mathbb{P}(H_{t}|\gamma)\mathrm{d}\mathcal{P_{\gamma}}$ satisfies $\int_{\left\{\gamma:\ a\leq a_{(\gamma)}\right\}}\mathbb{P}(H_{t}|\gamma)\mathrm{d}\mathcal{P_{\gamma}}=\int_{\left\{\gamma:\ a\geq a_{(\gamma)}\right\}}\mathbb{P}(H_{t}|\gamma)\mathrm{d}\mathcal{P_{\gamma}}$. Divide both sides of this equation by $\int_{\gamma}\mathbb{P}(H_{t}|\gamma)\mathrm{d}\mathcal{P_{\gamma}}$, we can conclude that
$\int_{\left\{\gamma:\ a\leq a_{(\gamma)}\right\}}\mathrm{d}\mathcal{P}\left({\gamma}|H_{t}\right)=\int_{\left\{\gamma:\ a\geq a_{(\gamma)}\right\}}\mathrm{d}\mathcal{P}\left({\gamma}|H_{t}\right),$
where $\mathcal{P}\left(\gamma|H_{t}\right)$ is the posterior distribution of $\gamma$. Hence $\texttt{Alg}^{*}(H)$ is the posterior median.
D.5 Proof of Proposition 4.3
Proof.
(1). For the linear bandits problem, we consider the following toy model. Consider two environments $\left\{\gamma_{1},\gamma_{2}\right\}$, the prior distribution $\mathcal{P}_{\gamma}$ is given by $\mathcal{P}(\gamma_{1})=\mathcal{P}(\gamma_{2})=\frac{1}{2}$. There is no context, the action space is 2-dimensional. Condition on $\gamma,a_{1},o_{1},\ldots,a_{t}$, the conditional distribution of $o_{t}$ (which contains the reward $R_{t}$ only) is the normal distribution $N\left(w_{\gamma}^{T}a_{t},1\right)$, where $\omega_{\gamma}$ will be given later.
Given $\gamma$ and $H_{t}$, $a_{t}$ is obtained by the posterior averaging
$a_{t}=\texttt{Alg}^{*}(H_{t})=\operatorname*{arg\,min}_{a\in\mathcal{A}}\int_{\gamma}||a-a^{*}(H_{t},\gamma)||^{2}\mathbb{P}(H_{t}|\gamma)\mathrm{d}\mathcal{P_{\gamma}}.$
Denote $a^{*}(H_{t},\gamma)=\mathop{\operatorname*{arg\,max}}\limits_{a\in\mathcal{A}}\ w_{\gamma}^{T}a$ by $a_{(\gamma)}$. Let $w_{\gamma_{1}}=(1,0)^{T}$, $w_{\gamma_{2}}=(0,1)^{T}$, and $\mathcal{A}=[-1,1]\times[-1,1]$, $a_{(\gamma_{1})}=(1,0)^{T}$, $a_{(\gamma_{2})}=(0,1)^{T}$.
Observe that by part (b) in Example 4.2 (see also D.4 for more details), we have
$a_{t}=\texttt{Alg}^{*}(H_{t})=\mathcal{P}\left(\gamma_{1}|H_{t}\right)\cdot a_{(\gamma_{1})}+\mathcal{P}\left(\gamma_{2}|H_{t}\right)\cdot a_{(\gamma_{2})},$
where $\mathcal{P}\left(\gamma|H_{t}\right)$ is the posterior of $\gamma$ condition on $H_{t}$, and is given by
$\mathcal{P}\left(\gamma_{i}|H_{t}\right)=\frac{\prod_{\tau=1}^{t-1}\mathbb{P}_{i}(o_{\tau}|a_{\tau})}{\sum_{i=1}^{2}\prod_{\tau=1}^{t-1}\mathbb{P}_{i}(o_{\tau}|a_{\tau})},\quad i=1,2.$
We now use induction to show $\mathcal{P}\left(\gamma_{i}|H_{t}\right)=\frac{1}{2}$ for any $t$, thereby having $a_{t}=\frac{1}{2}a_{(\gamma_{1})}+\frac{1}{2}a_{(\gamma_{2})}=\left(\frac{1}{2},\frac{1}{2}\right)^{T}$ does not change with respect to $t$.
Step 1. Since $\mathcal{P}\left(\gamma_{i}|H_{0}\right)=\mathcal{P}\left(\gamma_{i}\right)=\frac{1}{2}$, we have $a_{1}=\frac{1}{2}a_{(\gamma_{1})}+\frac{1}{2}a_{(\gamma_{2})}=(\frac{1}{2},\frac{1}{2})^{\top}$, and the conclusion holds for $t=1$.
Step 2. Now assume the conclusion holds for $t$, i.e., $\mathcal{P}\left(\gamma_{i}|H_{t}\right)=\frac{1}{2}$, and $a_{t}=\left(\frac{1}{2},\frac{1}{2}\right)^{T}$. Since $w_{\gamma_{1}}^{T}a_{t}=w_{\gamma_{2}}^{T}a_{t}=\frac{1}{2}$, we have $\mathbb{P}_{i}(o_{t}|a_{t})=\frac{1}{\sqrt{2\pi}}\exp\left(-\frac{\left(o_{t}-\frac{1}{2}\right)^{2}}{2}\right)$. Observe that
$\mathcal{P}(\gamma_{i}|H_{t+1})=\frac{\mathcal{P}(\gamma_{i},H_{t+1})}{\sum_{i^{\prime}=1}^{2}\mathcal{P}(\gamma_{i^{\prime}},H_{t+1})}=\frac{\mathcal{P}(H_{t})\mathcal{P}(\gamma_{i}|H_{t})\mathbb{P}_{i}(o_{t}|a_{t})}{\sum_{i^{\prime}=1}^{2}\mathcal{P}(H_{t})\mathcal{P}(\gamma_{i^{\prime}}|H_{t})\mathbb{P}_{i^{\prime}}(o_{t}|a_{t})}=\frac{\mathbb{P}_{i}(o_{t}|a_{t})}{\sum_{i^{\prime}=1}^{2}\mathbb{P}_{i^{\prime}}(o_{t}|a_{t})}=\frac{1}{2},$
which implies $a_{t+1}=\left(\frac{1}{2},\frac{1}{2}\right)^{T}$, and the conclusion holds for $t+1$.
Therefore, by the principle of induction, the conclusion holds. Then the regret for $\gamma$ is
$\text{Regret}(\texttt{Alg}^{*};\gamma)=\mathbb{E}\left[\sum_{t=1}^{T}r(X_{t},a_{t}^{*})-r(X_{t},a_{t})\right]=\frac{1}{2}T,$
which completes the first part of the proof.
(2). For the dynamic pricing problem, we consider the following toy model. Consider two environments $\left\{\gamma_{1},\gamma_{2}\right\}$, the prior distribution $\mathcal{P}_{\gamma}$ is given by $\mathcal{P}(\gamma_{1})=\mathcal{P}(\gamma_{2})=\frac{1}{2}$. There is no context, and the action space is 1-dimensional. The demand functions of them are set to be $o_{t}=2-a_{t}+\epsilon_{t}$ and $o_{t}=\frac{4}{5}-\frac{1}{5}\cdot a_{t}+\epsilon_{t}$ respectively, where $\epsilon_{t}\overset{i.i.d.}{\sim}\mathcal{N}(0,1)$. Accordingly, the optimal actions (which follow the same notations as in (1)) are then $a_{(\gamma_{1})}=1$ and $a_{(\gamma_{2})}=2$.
Given $\gamma$ and $H_{t}$, $a_{t}$ is achieved by the posterior averaging:
$a_{t}=\texttt{Alg}^{*}(H_{t})=\sum_{i=1}^{2}\mathcal{P}(\gamma_{i}|H_{t})\cdot a_{(\gamma_{i})}$
Condition on $\gamma,a_{1},o_{1},\ldots,a_{t}$, the conditional distribution of the demand $o_{t}$ is the normal distribution $\mathcal{N}\left(2-a_{t},1\right)$ for $\gamma_{1}$ and $\mathcal{N}\left(\frac{4-a_{t}}{5},1\right)$ for $\gamma_{2}$. Observe that $\tilde{a}=\frac{1}{2}a_{(\gamma_{1})}+\frac{1}{2}a_{(\gamma_{2})}=\frac{3}{2}$ satisfies that $2-\tilde{a}=\frac{4-\tilde{a}}{5}$ and $a_{1}=\tilde{a}$, then by the similar procedure in (1) we can conclude that $\mathcal{P}\left(\gamma_{i}|H_{t}\right)=\frac{1}{2}$ for any $t$. Therefore, $a_{t}=\frac{1}{2}a_{(\gamma_{1})}+\frac{1}{2}a_{(\gamma_{2})}=\frac{3}{2}$ does not change with respect to $t$. Therefore the regret for $\gamma$ is
$\text{Regret}(\texttt{Alg}^{*};\gamma)=\mathbb{E}\left[\sum_{t=1}^{T}r(X_{t},a_{t}^{*})-r(X_{t},a_{t})\right]=\frac{1}{4}T\cdot\mathbb{I}\left\{\gamma=\gamma_{1}\right\}+\frac{1}{20}T\cdot\mathbb{I}\left\{\gamma=\gamma_{2}\right\},$
which completes the second part of the proof.
D.6 Proof of Proposition 4.4
In the following proof, we assume Assumption B.1 holds. Recall that the posterior condition on history $H_{t}$ is $\mathcal{P}(\gamma_{i}|H_{t})=\frac{\prod_{{\tau}=1}^{t-1}\mathbb{P}_{\gamma_{i}}(o_{\tau}|H_{\tau},\texttt{TF}_{\hat{\theta}})}{\sum_{i=1}^{k}\prod_{{\tau}=1}^{t-1}\mathbb{P}_{\gamma_{i}}(o_{\tau}|H_{\tau},\texttt{TF}_{\hat{\theta}})}$, then for $\gamma\neq\gamma_{i}$, the log-likelihood ratio is
$\log\left(\frac{\mathcal{P}(\gamma|H_{t})}{\mathcal{P}(\gamma_{i}|H_{t})}\right)=\sum_{{\tau}=1}^{t-1}\log\left(\frac{\mathbb{P}_{\gamma}(o_{\tau}|H_{\tau},\texttt{TF}_{\hat{\theta}})}{\mathbb{P}_{\gamma_{i}}(o_{\tau}|H_{\tau},\texttt{TF}_{\hat{\theta}})}\right).$
Since $\log\left(\frac{\mathbb{P}_{\gamma}(o_{\tau}|H_{\tau},\texttt{TF}_{\hat{\theta}})}{\mathbb{P}_{\gamma_{i}}(o_{\tau}|H_{\tau},\texttt{TF}_{\hat{\theta}})}\right)$ is $\sigma^{2}$-subgaussian condition on $H_{t}$ with respect to environment $\gamma$, and $\left\{\log\left(\frac{\mathbb{P}_{\gamma}(o_{\tau}|H_{\tau},\texttt{TF}_{\hat{\theta}})}{\mathbb{P}_{\gamma_{i}}(o_{\tau}|H_{\tau},\texttt{TF}_{\hat{\theta}})}\right)-\mathbb{E}\left[\log\left(\frac{\mathbb{P}_{\gamma}(o_{\tau}|H_{\tau},\texttt{TF}_{\hat{\theta}})}{\mathbb{P}_{\gamma_{i}}(o_{\tau}|H_{\tau},\texttt{TF}_{\hat{\theta}})}\right)\Big{|}H_{\tau}\right]\right\}_{{\tau}=1}^{t}$ is a martingale difference sequence, by the concentration bounds for martingale difference sequence (Section 2.2.2 in [Wainwright, 2019]), with probability at least $1-h$, we have
$\log\left(\frac{\mathcal{P}(\gamma|H_{t})}{\mathcal{P}(\gamma_{i}|H_{t})}\right)$ $\geq\sum_{{\tau}=1}^{t-1}\mathbb{E}\left[\log\left(\frac{\mathbb{P}_{\gamma}(o_{\tau}|H_{\tau},\texttt{TF}_{\hat{\theta}})}{\mathbb{P}_{\gamma_{i}}(o_{\tau}|H_{\tau},\texttt{TF}_{\hat{\theta}})}\right)\Big{|}H_{\tau}\right]-\sqrt{2\sigma^{2}(t-1)\log\left(\frac{1}{h}\right)}$
$\geq(t-1)\Delta-\sqrt{2\sigma^{2}(t-1)\log\left(\frac{1}{h}\right)}.$
[…]
$\log\left(\frac{\mathcal{P}(\gamma|H_{t})}{\mathcal{P}(\gamma_{i}|H_{t})}\right)$ $\geq(t-1)\Delta-\sqrt{2\sigma^{2}(t-1)\log\left(\frac{k}{h}\right)},$
which implies that
$\mathcal{P}(\gamma|H_{t})\geq\exp\left((t-1)\Delta-\sqrt{2\sigma^{2}(t-1)\log\left(\frac{k}{h}\right)}\right)\mathcal{P}(\gamma_{i}|H_{t}).$
In view of this, we have
$0\leq\sum_{\gamma_{i}\neq\gamma}\mathcal{P}(\gamma_{i}|H_{t})\leq\frac{k-1}{k-1+\exp\left((t-1)\Delta-\sqrt{2\sigma^{2}(t-1)\log\left(\frac{k}{h}\right)}\right)}.$
Notice that the term $k-1$ in the denominator, derived from the condition $\mathop{\sum}\limits_{i=1}^{k}\mathcal{P}(\gamma_{i}|H_{t})=1$, is essential for getting the final regret bound.
Now we start to bound the regret. Recall that under the square loss, $\texttt{Alg}^{*}(H_{t})=\mathop{\sum}\limits_{i=1}^{k}a^{*}(H_{t},\gamma_{i})\mathcal{P}(\gamma_{i}|H_{t})$, where $a^{*}(H_{t},\gamma_{i})$ is the optimal action in environment $\gamma_{i}$. Then conditioned on $H_{t}$, with respect to $\gamma$ we have
$\mathbb{E}\left[r(X_{t},a_{t}^{*})-r(X_{t},\texttt{TF}_{\hat{\theta}}(H_{t}))|H_{t}\right]$
$=$ $\mathbb{E}\left[r(X_{t},a^{*}(H_{t},\gamma))-r\left(X_{t},\texttt{Alg}^{*}(H_{t})\right)|H_{t}\right]+\mathbb{E}\left[r\left(X_{t},\texttt{Alg}^{*}(H_{t})\right)-r\left(X_{t},\texttt{TF}_{\hat{\theta}}(H_{t})\right)|H_{t}\right].$
Recall that from Proposition B.2, the regret satisfies the surrogate property with respect to squared loss
$r(X_{t},a^{*}(H_{t},\gamma))-r\left(X_{t},\texttt{Alg}^{*}(H_{t})\right)\leq C||a^{*}(H_{t},\gamma)-\texttt{Alg}^{*}(H_{t})||_{2}^{2},$
this yields
$\mathbb{E}\left[r(X_{t},a_{t}^{*})-r(X_{t},\texttt{TF}_{\hat{\theta}}(H_{t}))|H_{t}\right]$
$\leq$ $C\mathbb{E}\left[||a^{*}(H_{t},\gamma)-\texttt{Alg}^{*}(H_{t})||_{2}^{2}\mid H_{t}\right]+\epsilon$
$=$ $C\mathbb{E}\left[\Big{|}\Big{|}\sum_{\gamma_{i}\neq\gamma}^{m}\left(a^{*}(H_{t},\gamma)-a^{*}(H_{t},\gamma_{i})\right)\mathcal{P}(\gamma_{i}|H_{t})\Big{|}\Big{|}_{2}^{2}\Big{|}H_{t}\right]+\epsilon$
$\leq$ $CD^{2}h+CD^{2}(1-h)\frac{k-1}{k-1+\exp\left((t-1)\Delta-\sqrt{2\sigma^{2}(t-1)\log\left(\frac{k}{h}\right)}\right)}+\epsilon,$
where $D$ is the diameter of the action space. Hence we have
$\mathrm{Regret}(\texttt{TF}_{\hat{\theta}};\gamma)=$ $\mathbb{E}\left[\sum_{t=1}^{T}r(X_{t},a^{*}(H_{t},\gamma))-r\left(X_{t},\texttt{TF}_{\hat{\theta}}(H_{t})\right)\right]$
$\leq$ $CD^{2}\left(hT+\sum_{t=1}^{T}\frac{k-1}{k-1+\exp\left((t-1)\Delta-\sqrt{2\sigma^{2}(t-1)\log\left(\frac{k}{h}\right)}\right)}\right)+\epsilon T$
$\leq$ $O\left((\epsilon+h)T+\frac{\log k}{\Delta}-\frac{\log h}{\Delta}\right).$
Taking $h=\epsilon$ completes the proof.
D.6.1 Justification for the conditions in Proposition 4.4
From now on, we provide some justification for the conditions in Proposition 4.4.
(1). We illustrate two examples (linear badits and dynamic pricing) to show the validity of the condition that the conditional log-likelihood ratio $\log\left(\frac{\mathbb{P}_{\gamma_{i}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})}{\mathbb{P}_{\gamma_{j}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})}\right)$ is $\sigma^{2}$-subgaussian with respect to $\gamma_{i}$.
  • For linear bandits problem, given $\gamma,a_{1},o_{1},\ldots,a_{t}$, assume the conditional distribution of $o_{t}$ is the normal distribution $N\left(w_{\gamma}^{T}\texttt{TF}_{\hat{\theta}}(H_{t}),\tilde{\sigma}^{2}\right)$, then we have
    $\log\left(\frac{\mathbb{P}_{\gamma_{i}}(o_{t}|H_{t},\texttt{TF}_{\hat{\theta}})}{\mathbb{P}_{\gamma_{j}}\left(o_{t}|H_{t},\texttt{TF}_{\hat{\theta}}\right)}\right)$
    $=$ $\frac{\left(o_{t}-w_{\gamma_{j}}^{T}\texttt{TF}_{\hat{\theta}}(H_{t})\right)^{2}}{2\tilde{\sigma}^{2}}-\frac{\left(o_{t}-w_{\gamma_{i}}^{T}\texttt{TF}_{\hat{\theta}}(H_{t})\right)^{2}}{2\tilde{\sigma}^{2}}$
    $=$ $\frac{\left(w_{\gamma_{i}}^{T}\texttt{TF}_{\hat{\theta}}(H_{t})-w_{\gamma_{j}}^{T}\texttt{TF}_{\hat{\theta}}(H_{t})\right)o_{t}}{\tilde{\sigma}^{2}}+\frac{\left(w_{\gamma_{j}}^{T}\texttt{TF}_{\hat{\theta}}(H_{t})\right)^{2}-\left(w_{\gamma_{i}}^{T}\texttt{TF}_{\hat{\theta}}(H_{t})\right)^{2}}{2\tilde{\sigma}^{2}},$
    which is $D^{2}$-subgaussian with respect to $\gamma_{1}$, where $D$ is the bounded parameter of observation space and action space.
  • For dynamic pricing problem, given $\gamma,a_{1},o_{1},\ldots,a_{t}$, assume the conditional distribution of the demand $o_{t}$ is the normal distribution $\mathcal{N}\left(e_{\gamma}-f_{\gamma}a_{t},1\right)$, where $(e_{\gamma},f_{\gamma})$ is in the ball with center 0 and radius $\frac{D}{2}$, then we have
    $\log\left(\frac{\mathbb{P}_{\gamma_{i}}(o_{t}|H_{t},\texttt{TF}_{\hat{\theta}})}{\mathbb{P}_{\gamma_{j}}\left(o_{t}|H_{t},\texttt{TF}_{\hat{\theta}}\right)}\right)$
    $=$ $\frac{\left(o_{t}-\left(e_{\gamma_{j}}-f_{\gamma_{j}}\texttt{TF}_{\hat{\theta}}(H_{t})\right)\right)^{2}}{2\tilde{\sigma}^{2}}-\frac{\left(o_{t}-\left(e_{\gamma_{i}}-f_{\gamma_{1}}\texttt{TF}_{\hat{\theta}}(H_{t})\right)\right)^{2}}{2\tilde{\sigma}^{2}}$
    $=$ $\frac{\left((f_{\gamma_{j}}-f_{\gamma_{i}})\texttt{TF}_{\hat{\theta}}(H_{t})+e_{\gamma_{i}}-e_{\gamma_{j}}\right)o_{\tau}}{\tilde{\sigma}^{2}}+\frac{\left(e_{\gamma_{j}}-f_{\gamma_{j}}\texttt{TF}_{\hat{\theta}}(H_{t})\right)^{2}-\left(e_{\gamma_{i}}-f_{\gamma_{i}}\texttt{TF}_{\hat{\theta}}(H_{t})\right)^{2}}{2\tilde{\sigma}^{2}},$
    which is $D^{2}+D$-subgaussian with respect to $\gamma_{1}$.
(2). To validate the condition that $\text{KL}\left(\mathcal{P}_{\gamma_{i}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})\|\mathcal{P}_{\gamma_{j}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})\right)\geq\Delta$, we illustrate with two examples: linear bandits and dynamic pricing. In these setting, we show that there exists a region in the neighborhood of $\texttt{Alg}^{*}(H_{t})$ where $\texttt{TF}_{\hat{\theta}}(H_{t})$ will satisfy this condition.
More specifically we show that this condition can be achieved when $\texttt{TF}_{\hat{\theta}}(H_{t})$ is defined as
$\texttt{TF}_{\hat{\theta}}(H_{t})=\left\{\begin{array}[]{lcl}\texttt{Alg}^{*}(H_{t}),&&\text{with probability $1-\epsilon$},\\ \bar{a},&&{\text{with probability $\epsilon$}},\\ \end{array}\right.$
for some $\bar{a}$.
  • For linear bandits problem,
    $\text{KL}\left(\mathcal{P}_{\gamma_{i}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})\|\mathcal{P}_{\gamma_{j}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})\right)$
    $=$ $\epsilon\cdot\frac{\left(w_{\gamma_{i}}^{T}\bar{a}-w_{\gamma_{j}}^{T}\bar{a}\right)^{2}}{2\tilde{\sigma}^{2}}+(1-\epsilon)\cdot\frac{\left(w_{\gamma_{i}}^{T}\texttt{Alg}^{*}(H_{\tau})-w_{\gamma_{j}}^{T}\texttt{Alg}^{*}(H_{\tau})\right)^{2}}{2\tilde{\sigma}^{2}}.$
    Since the second term is always non-negative, one can choose some $\bar{a}$ which only depends on $w_{\gamma_{1}},\ldots,w_{\gamma_{k}}$ such that $\text{KL}\left(\mathcal{P}_{\gamma_{i}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})\|\mathcal{P}_{\gamma_{j}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})\right)\geq\epsilon\tilde{C}$ for some constant $\tilde{C}$, which also shows the trade-off between $\Delta$ and $\epsilon$ in Proposition 4.4.
  • For dynamic pricing problem,
    $\text{KL}\left(\mathcal{P}_{\gamma_{i}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})\|\mathcal{P}_{\gamma_{j}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})\right)$
    $=$ $\epsilon\cdot\frac{\left((f_{\gamma_{j}}-f_{\gamma_{i}})\bar{a}+e_{\gamma_{i}}-e_{\gamma_{j}}\right)^{2}}{2\tilde{\sigma}^{2}}+(1-\epsilon)\cdot\frac{\left((f_{\gamma_{j}}-f_{\gamma_{i}})\texttt{Alg}^{*}(H_{\tau})+e_{\gamma_{i}}-e_{\gamma_{j}}\right)^{2}}{2\tilde{\sigma}^{2}}.$
    Then we can choose some $\bar{a}$ which only depends on $(e_{\gamma_{1}},f_{\gamma_{1}}),\ldots,(e_{\gamma_{k}},f_{\gamma_{k}})$ such that $\text{KL}\left(\mathcal{P}_{\gamma_{i}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})\|\mathcal{P}_{\gamma_{j}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})\right)\geq\epsilon\tilde{C}\triangleq\Delta$ for some constant $\tilde{C}$.
Appendix E Details for Numerical Experiments
E.1 Transformer Architecture and Algorithm 1 Implementation
E.1.1 Transformer for sequential decision making
We adopt the transformer architecture for regression tasks from [Garg et al., 2022] to solve sequential decision-making tasks, with modifications tailored to our specific setting.
Prompt. At time $t$, the prompt consists of two types of elements derived from the history $H_{t}$: (i) the “feature” elements, which stack the context $X_{\tau}\in\mathbb{R}^{d}$ at each time $\tau\leq t$ and the observation $o_{\tau-1}\in\mathbb{R}^{k}$ from $\tau-1$ (with $o_{0}$ set as a $k$-dimensional zero vector in our experiments), i.e., $\{(X_{\tau},o_{\tau-1})\}_{\tau=1}^{t}\subset\mathbb{R}^{d+k}$. These elements serve as the “features” in the prediction; and (ii) the “label” elements, which are the actions $\{a_{\tau}\}_{\tau=1}^{t-1}$ in $H_{t}$. Therefore, the prompt contains $12t-$ 1 elements in total.
Architecture. The transformer is based on the GPT-2 family [Radford et al., 2018]. It uses two learnable linear transformations to map each “feature” and “label” element into vectors in the embedding space respectively (similar to tokens in language models). These vectors are processed through the GPT-2’s attention mechanism, resulting in a vector that encapsulates relevant contextual information. This vector then undergoes another learnable linear transformation, moving from the embedding space to the action $\mathcal{A}$, ultimately resulting in the prediction of $a^{*}_{t}$. In our experiments, the GPT-2 model has 12 layers, 16 attention heads, and a 256-dimensional embedding space. The experiments are conducted on 2 A100 GPUs with DistributedDataParallel method of Pytorch.
E.1.2 Algorithm 1 implementation and pretraining details
During pre-training, we use the AdamW optimizer with a learning rate of $10^{-4}$ and a weight decay of $10^{-4}$. The dropout rate is set to 0.05. For implementing Algorithm 1 in our experiments, we set the total iterations $M=130$ with an early training phase of $M_{0}=50$. The number of training sequences per iteration is $n=1500\times 64$, which are randomly split into 1500 batches with a batch size of 64. The transformer’s parameter $\theta$ is optimized to minimize the averaged loss of each batch. As suggested in Proposition B.2, we select the cross-entropy loss for the multi-armed bandits, squared loss for the dynamic pricing, and absolute loss for the linear bandits and newsvendor problem.
To reduce computation costs from sampling histories during the early training phase, instead of sampling a new dataset $D_{m}$ in each iteration (as described in Algorithm 1), we initially sample $10^{6}$ data samples before the pre-training phase as an approximation of $\mathcal{P}_{\gamma,f}$. Each sample in the batch is uniformly sampled from these $10^{6}$ data samples.
To reduce transformer inference costs during the mixed training phase, we reset the number of training sequences to $n=15\times 64$ per iteration and set the ratio $\kappa=1/3$, meaning $10\times 64$ samples are from $\mathcal{P}_{\gamma,\texttt{TF}_{\theta_{m}}}$ and $5\times 64$ samples are from $\mathcal{P}_{\gamma,f}$ (sampled from the pool of pre-generated $10^{6}$ samples as before). Consequently, we adjust the number of batches from 1500 to 50 to fit the smaller size of training samples, keeping the batch size at 64.
Decision Function $f$ in Pre-Training. To mitigate the OOD issue mentioned in Section 3, we aim for the decision function $f$ to approximate $\texttt{Alg}^{*}$ or $\texttt{TF}_{\hat{\theta}}$. However, due to the high computation cost of $\texttt{Alg}^{*}$ and the unavailability of $\texttt{TF}_{\hat{\theta}}$, we set $f(H_{t})=a^{*}_{t}+\epsilon^{\prime}_{t}$, where $\epsilon^{\prime}_{t}$ is random noise simulating the suboptimality of $\texttt{Alg}^{*}$ or $\texttt{TF}_{\hat{\theta}}$ and independent of $H_{t}$. As we expect such suboptimality to decrease across $t$, we also reduce the influence of $\epsilon^{\prime}_{t}$ across $t$. Thus, in our experiments, $\epsilon^{\prime}_{t}$ is defined as:
$\epsilon^{\prime}_{t}\sim\begin{cases}0&\text{w.p.}\ \max\{0,1-\frac{2}{\sqrt{t}}\},\\ \text{Unif}[-1,1]&\text{w.p.}\ \min\{1,\frac{2}{\sqrt{t}}\},\end{cases}$
where for the multi-armed bandits we replace $\text{Unif}[-1,1]$ by $\text{Unif}\{-2,-1,1,2\}$ and for the linear bandits with dimension $d>1$ we apply the uniform distribution of the set $[-1,-1]^{d}$. We further project $f(H_{t})$ into $\mathcal{A}$ when $f(H_{t})\notin\mathcal{A}$.
Curriculum. Inspired by Garg et al. [2022] in the regression task, we apply curriculum training to potentially speed up Algorithm 1. This technique uses “simple” task data initially and gradually increases task complexity during the training. Specifically, we train the transformer on samples with a smaller horizon $\tilde{T}=20$ (generated by truncating samples from $\mathcal{P}_{\gamma,f}$ or $\mathcal{P}_{\gamma,\texttt{TF}_{\theta_{m}}}$) at the beginning and gradually increase the sample horizon to the target $T=100$. We apply this curriculum in both the early training phase ($m\leq 50$) and the mixed training phase ($m>50$). The last 30 iterations focus on training with $\tilde{T}=100$ using non-truncated samples from $\mathcal{P}_{\gamma,\texttt{TF}_{\theta}}$ to let the transformer fit more on the non-truncated samples. The exact setting of $\tilde{T}$ is as follows:
$\tilde{T}=\begin{cases}20\times(m\%10+1)&\text{when}\ m=1,\ldots,50,\\ 20\times(m\%10-4)&\text{when}\ m=51,\ldots,100,\\ 100&\text{when}\ m=100,\ldots,130.\end{cases}$
E.2 Environment Generation
Throughout the paper, we set the horizon $T=100$ for all tasks. The environment distribution $\mathcal{P}_{\gamma}$ is defined as follows for each task:
[…]
  • Linear bandits: We consider the dimension of actions/arms $d=2$ and $\mathcal{A}=\mathcal{B}(0,1)$, i.e., a unit ball centered at the origin (with respect to the Euclidean norm). The environment parameter $\gamma$ encapsulates the (linear) reward function’s parameter $w\in\{\tilde{w}\in\mathbb{R}^{2}:\|\tilde{w}\|_{2}=1\}$, where the reward of arm $a$ is sampled from $\mathcal{N}(w^{\top}a,0.2)$. Thus, the environment distribution $\mathcal{P}_{\gamma}$ is defined by the distribution of $w$. We set $w$ to be uniformly sampled from the unit sphere, and the corresponding optimal arm being $a^{*}=w$.
[…]
Finite Pool of Environments. To study the behavior and test the performance of the transformer on finite possible environments (e.g., to see if and how $\texttt{TF}_{\hat{\theta}}$ converges to $\texttt{Alg}^{*}$), we also consider a finite pool of environments as the candidates of $\gamma$. Specifically, for some tasks we first sample finite environments (e.g., $\gamma_{1},\ldots,\gamma_{4}$) i.i.d. following the sampling rules previously described. We then set the environment distribution $\mathcal{P}_{\gamma}$ as a uniform distribution over the pool of sampled environments (e.g., $\{\gamma_{1},\ldots,\gamma_{4}\}$). We should notice this pool of environments does not restrict the context generation (if any): in both the pretraining and testing, contexts are generated following the rules previously described and are independent of the given finite pool.
Tasks with Two Demand Types. To test the transformer’s performance on tasks with a mixture of different demand types, we also consider the square demand function family for the dynamic pricing and newsvendor tasks. Specifically, besides the linear demand function family, we also consider $d^{\text{sq}}(X_{t},a_{t})=(w_{1}^{\top}X_{t})^{2}-(w_{2}^{\top}X_{t})\cdot a_{t}+\epsilon_{t}$ in dynamic pricing and $d^{\text{sq}}(X_{t},a_{t})=(w^{\top}X_{t})^{2}+\epsilon_{t}$ in the newsvendor problem. Thus, with a slight abuse of notation, we can augment $\gamma$ such that it parameterizes the type of demand (linear or square). In experiments related to multi-type demands, the probability of each type being sampled is $1/2$, while the distributions over other parameters of the environment remain the same as in the linear demand case.
E.3 Benchmark Algorithms
E.3.1 Multi-armed bandits
  • Upper Confidence Bound (UCB) [Lattimore and Szepesvári, 2020]: Given $H_{t}$, the action $a_{t}$ is chosen by $a_{t}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\hat{r}_{a}+\frac{\sqrt{2\log T}}{\min\{1,n_{a}\}}$, where $n_{a}$ is the number of pulling times of arm $a$ before time $t$ and $\hat{r}_{a}$ is the empirical mean reward of $a$ based on $H_{t}$.
  • Thomspon sampling (TS) [Russo et al., 2018]: Given $H_{t}$, the action $a_{t}$ is chosen by $a_{t}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\tilde{r}_{a}$, where each $\tilde{r}_{a}\sim\mathcal{N}\left(\hat{r}_{a},\frac{\sqrt{2\log T}}{\min\{1,n_{a}\}}\right)$ is the sampled reward of arm $a$ and $\hat{r}_{a},n_{a}$ follow the same definitions as in UCB.
  • $\texttt{Alg}^{*}$: Given $H_{t}$ from a task with a pool of $|\Gamma|$ environments $\{\gamma_{1},\ldots,\gamma_{|\Gamma|}\}$, the action $a_{t}$ is chosen by the posterior sampling defined in Algorithm 2. The posterior distribution can be computed by
    $\mathcal{P}(\gamma_{i}|H_{t})=\frac{\exp(-\frac{1}{\sigma^{2}}\sum_{\tau=1}^{t-1}(o_{\tau}-r^{i}_{a_{\tau}})^{2})}{\sum_{i^{\prime}=1}^{|\Gamma|}\exp(-\frac{1}{\sigma^{2}}\sum_{\tau=1}^{t-1}(o_{\tau}-r^{i^{\prime}}_{a_{\tau}})^{2})},$
    where $r^{i}_{a}$ is the expected reward of $a$ in environment $\gamma_{i}$ and $\sigma^{2}$ is the variance of the noise (which equals to $0.2$ in our experiments).
E.3.2 Linear bandits
  • LinUCB [Chu et al., 2011]: Given $H_{t}$, we define $\Sigma_{t}=\sum_{\tau=1}^{t-1}a_{\tau}a_{\tau}^{\top}+\sigma^{2}I_{d}$, where $\sigma^{2}$ is the variance of the reward noise. The action $a_{t}$ is chosen by $a_{t}\in\operatorname*{arg\,max}_{a\in\mathcal{A}}\hat{w}_{t}^{\top}a+\sqrt{2\log T}\|a\|_{\Sigma_{t}^{-1}}$, where $\hat{w}_{t}=\Sigma_{t}^{-1}(\sum_{\tau=1}^{t-1}o_{\tau}\cdot a_{\tau})$ is the current estimation of $w$ from $H_{t}$.
  • LinTS [Agrawal and Goyal, 2013]: Given $H_{t}$, the action $a_{t}$ is chosen by $a_{t}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\tilde{w}_{t}^{\top}a$, where $\tilde{w}_{t}\sim\mathcal{N}\left(\hat{w}_{t},\sqrt{2\log T}\Sigma_{t}^{-1}\right)$ as the sampled version of $w$ and $\hat{w}_{t},\Sigma_{t}$ follow the same definitions in LinUCB.
  • $\texttt{Alg}^{*}$: Given $H_{t}$ from a task with a pool of $|\Gamma|$ environments $\{\gamma_{1},\ldots,\gamma_{|\Gamma|}\}$, the action $a_{t}$ is chosen by the posterior median defined in Algorithm 2. The posterior distribution can be computed by
    $\mathcal{P}(\gamma_{i}|H_{t})=\frac{\exp(-\frac{1}{\sigma^{2}}\sum_{\tau=1}^{t-1}(o_{\tau}-w_{i}^{\top}a_{\tau})^{2})}{\sum_{i^{\prime}=1}^{|\Gamma|}\exp(-\frac{1}{\sigma^{2}}\sum_{\tau=1}^{t-1}(o_{\tau}-w^{\top}_{i^{\prime}}a_{\tau})^{2})},$
    where $w_{i}$ is the reward function parameter in environment $\gamma_{i}$ and $\sigma^{2}$ is the variance of the noise (which equals to $0.2$ in our experiments).
E.3.3 Dynamic pricing
All the benchmark algorithms presented below assume the demand model belongs to the linear demand function family as defined in Appendix E.2 (which can be mis-specified when we deal with dynamic pricing problems with two demand types).
  • Iterative least square estimation (ILSE) [Qiang and Bayati, 2016]: At each $t$, it first estimates the unknown demand parameters $(w_{1},w_{2})$ by applying a ridge regression based on $H_{t}$, and chooses $a_{t}$ as the optimal action according to the estimated parameters and the context $X_{t}$. Specifically, we denote $(\hat{w}_{1,t},-\hat{w}_{2,t})=\Sigma_{t}^{-1}(\sum_{\tau=1}^{t-1}o_{\tau}\cdot z_{\tau})$ as the estimation of $(w_{1},-w_{2})$ through a ridge regression, where $z_{\tau}=(X_{\tau},a_{\tau}\cdot X_{\tau})$ serves as the “feature vector” and $\Sigma_{t}=\sum_{\tau=1}^{t-1}z_{\tau}z_{\tau}^{\top}+\sigma^{2}I_{d}$ ($\sigma^{2}$ is the variance of the demand noise). Then the action is chosen as the optimal one by treating $(\hat{w}_{1,t},\hat{w}_{2,t})$ as the true parameter: $a_{t}=\frac{\hat{w}_{1,t}^{\top}X_{t}}{2\hat{w}_{2,t}^{\top}X_{t}}$.
  • Constrained iterated least squares (CILS) [Keskin and Zeevi, 2014]: It is similar to the ILSE algorithm except for the potential explorations when pricing. Specifically, we follow the notations in ILSE and denote $\hat{a}_{t}=\frac{\hat{w}_{1,t}^{\top}X_{t}}{2\hat{w}_{2,t}^{\top}X_{t}}$ as the optimal action by treating $(\hat{w}_{1,t},\hat{w}_{2,t})$ as the true parameter (the chosen action of ILSE), and denote $\bar{a}_{t-1}$ as the empirical average of the chosen actions so far. Then the CILS chooses
    $a_{t}=\begin{cases}\bar{a}_{t-1}+\text{sgn}(\delta_{t})\frac{t^{-\frac{1}{4}}}{10},&\text{if }|\delta_{t}|<\frac{1}{10}t^{-\frac{1}{4}},\\ \hat{a}_{t},&\text{otherwise },\end{cases}$
    where $\delta_{t}=\hat{a}_{t}-\bar{a}_{t-1}$. The intuition is that if the tentative price $\hat{a}_{t}$ stays too close to the history average, it will introduce a small perturbation around the average as price experimentation to encourage parameter learning.
  • Thomspon sampling for pricing (TS) [Wang et al., 2021]: Like the ILSE algorithm, it also runs a regression to estimate $(w_{1},w_{2})$ based on the history data, while the chosen action is the optimal action of a sampled version of parameters. Specifically, we follow the notations in ILSE and then TS chooses
    $a_{t}=\frac{\tilde{\alpha}}{2\cdot\tilde{\beta}},$
    where $(\tilde{\alpha},\tilde{\beta})\sim\mathcal{N}((\hat{w}^{\top}_{1,t}X_{t},\hat{w}_{2,t}^{\top}X_{t}),\tilde{\Sigma}_{t}^{-1})\in\mathbb{R}^{2}$ are the sampled parameters of the “intercept” and “price coefficient” in the linear demand function and
    $\tilde{\Sigma}_{t}=\left(\begin{pmatrix}X_{t}&\bm{0}\\ \bm{0}&X_{t}\end{pmatrix}^{\top}\Sigma_{t}^{-1}\begin{pmatrix}X_{t}&\bm{0}\\ \bm{0}&X_{t}\end{pmatrix}\right)^{-1}$
    is the empirical covariance matrix given $X_{t}$.
  • $\texttt{Alg}^{*}$: Given $H_{t}$ from a task with a pool of $|\Gamma|$ environments $\{\gamma_{1},\ldots,\gamma_{|\Gamma|}\}$, the action $a_{t}$ is chosen by the posterior averaging defined in Algorithm 2. To compute the posterior distribution, we follow the notations in ILSE and denote $w=(w_{1},w_{2})$ as the stacked vector of parameters, then the posterior distribution is
    $\mathcal{P}(\gamma_{i}|H_{t})=\frac{\exp(-\frac{1}{\sigma^{2}}\sum_{\tau=1}^{t-1}(o_{\tau}-w_{i}^{\top}z_{\tau})^{2})}{\sum_{i^{\prime}=1}^{|\Gamma|}\exp(-\frac{1}{\sigma^{2}}\sum_{\tau=1}^{t-1}(o_{\tau}-w^{\top}_{i^{\prime}}z_{\tau})^{2})},$
    where $w_{i}$ is the demand function parameter in environment $\gamma_{i}$ and $\sigma^{2}$ is the variance of the noise (which equals to $0.2$ in our experiments).
E.3.4 Newsvendor
All the benchmark algorithms presented below assume the demand model belongs to the linear demand function family as defined in Appendix E.2 (which can be mis-specified when we deal with newsvendor problems with two demand types).
  • Empirical risk minimization (ERM) [Ban and Rudin, 2019]: Since the optimal action $a^{*}_{t}$ is the $\frac{1}{1+h}$ quantile of the random variable $w^{\top}X_{t}+\epsilon_{t}$ [Ding et al., 2024], ERM conducts a linear quantile regression based on the observed contexts and demands $\{(X_{\tau},o_{\tau})\}_{\tau=1}^{t-1}$ to predict the $\frac{1}{1+h}$ quantile on $X_{t}$.
  • Feature-based adaptive inventory algorithm (FAI) [Ding et al., 2024]: FAI is an online gradient descent style algorithm aiming to minimize the cost $\sum_{t=1}^{T}h\cdot(a_{t}-o_{t})^{+}+l\cdot(o_{t}-a_{t})^{+}$ (we set $l=1$ in our experiments). Specifically, it chooses $a_{t}=\tilde{w}_{t}^{\top}X_{t}$, where
    $\tilde{w}_{t}=\begin{cases}\tilde{w}_{t-1}-\frac{h}{\sqrt{t}}\cdot x_{t-1},&\text{if }o_{t-1}<a_{t-1},\\ \tilde{w}_{t-1}+\frac{l}{\sqrt{t}}\cdot x_{t-1},&\text{otherwise},\end{cases}$
    is the online gradient descent step and $\tilde{w}_{0}$ can be randomly sampled in $[0,1]^{d}$.
  • $\texttt{Alg}^{*}$: Given $H_{t}$ from a task with a pool of $|\Gamma|$ environments $\{\gamma_{1},\ldots,\gamma_{|\Gamma|}\}$, the action $a_{t}$ is chosen by the posterior median as shown in Algorithm 2. To compute the posterior distribution, we denote $\bar{\epsilon}_{\gamma}$ and $\beta_{\gamma}$ as the noise upper bound and demand function parameter of $\gamma$ at $\tau\leq t-1$, and define the event $\mathcal{E}_{\gamma,\tau}=\left\{0\leq o_{\tau}-\beta_{\gamma}^{\top}X_{\tau}\leq\bar{\epsilon}\right\}$ to indicate the feasibility of environment $\gamma$ from $(X_{\tau},o_{\tau})$, and denote ¯$\bar{\mathcal{E}}_{\gamma,t}=\bigcap_{\tau=1}^{\tau-1}\mathcal{E}_{\gamma,\tau}$ to indicate the feasibility at $t$. Then the posterior distribution of the underlying environment is
    $\mathcal{P}(\gamma_{i}|H_{t})=\frac{\mathbbm{1}_{\bar{\mathcal{E}}_{\gamma_{i},t}}\cdot\bar{\epsilon}_{\gamma_{i}}^{1-t}}{\sum_{i^{\prime}=1}^{|\Gamma|}\mathbbm{1}_{\bar{\mathcal{E}}_{\gamma_{i^{\prime}},t}}\cdot\bar{\epsilon}_{\gamma_{i^{\prime}}}^{1-t}}.$
E.4 Details for Figures
E.4.1 Details for Figure 1
Figure 1 (a).
In Figure 1 (a), we independently implement Algorithm 1 to train two transformer models on a dynamic pricing task, which has an infinite support of the prior environment distribution $\mathcal{P}_{\gamma}$ and $6$-dimensional contexts (more details can be found in Appendix E.2). For both models, the training parameters are identical except for the $M_{0}$ in Algorithm 1: the first curve (blue, thick line) uses $M_{0}=50$, while the second curve (orange, dashed line) uses $M_{0}=130$, meaning no transformer-generated data is utilized during training. All other parameters follow the configuration detailed in Appendix E.1. The figure shows the mean testing regret at each training iteration across 128 randomly sampled environments, with the shaded area representing the standard deviation. We further provide more results in Appendix C.1.
Figure 1 (b).
In Figure 1 (b), we consider a dynamic pricing task with a pool of $8$ linear demand functions and has $6$-dimensional contexts. We generate $30$ sequences from $\texttt{TF}_{\theta_{m}}$ across different training iterations for this task: before using the transformer-generated data in the training ($m=40,50$) and after using such data ($m=70,80,110,120$). All generated sequences share the same context sequence $\{X_{t}\}_{t=1}^{T}$ to control the effect of contexts on the chosen actions and the resulting observations. We stack the first $20$ actions and observations into a single sample point and use t-SNE method to visualize these high-dimension points as shown in Figure 1 (b). We further apply the same method to generate points from $\texttt{Alg}^{*}$ to study the behavior of TF.
We can observe that: (i) compared to the points from TF which have not been trained on self-generated data (i.e., $\texttt{TF}_{\theta_{40}},\texttt{TF}_{\theta_{50}}$), the points from the trained ones are closer to the $\texttt{Alg}^{*}$’s, i.e., the expected decision rule of a well-trained TF; (ii) When being trained with more self-generated data, the points from $\texttt{TF}_{\theta_{m}}$ get closer to the $\texttt{Alg}^{*}$’s. These observations suggest that utilizing the self-generated data (like in Algorithm 1) can help mitigate the OOD issue as discussed in Section 3: the data from $\texttt{TF}_{\theta_{m}}$ can get closer to the target samples from $\texttt{Alg}^{*}$ during the training and thus the pre-training data is closer to the testing data.
E.4.2 Details for Figure 2.
For Figure 2, subfigure (a) is based on a dynamic pricing task with $6$-dimensional contexts and (b) is based on a newsvendor task with $4$-dimensional contexts. Both tasks have a pool of $4$ environments, i.e., the support of $\mathcal{P}_{\gamma}$ only contains $4$ environments, with linear demand function type. Each subfigure is based on a sampled environment with a sampled sequence of contexts $\{X_{t}\}_{t=1}^{30}$ from the corresponding task. For each task, the data generation process follows the description detailed in Appendix E.2, and the architecture of $\texttt{TF}_{\hat{\theta}}$ and the definition of $\texttt{Alg}^{*}$ can be found in Appendix E.1 and Appendix E.3 respectively. We further provide more results and discussions in Appendix C.2.